### v1

# Comparing Bad Apples to Good Oranges: Aligning Large Language Models via Joint Preference Optimization

Abstract
A common technique for aligning large language models (LLMs) relies on acquiring human preferences by comparing multiple generations conditioned on a fixed context. This only leverages the pairwise comparisons when the generations are placed in an identical context. However, such conditional rankings often fail to capture the complex and multidimensional aspects of human preferences. In this work, we revisit the traditional paradigm of preference acquisition and propose a new axis that is based on eliciting preferences jointly over the instruction-response pairs. While prior preference optimizations are designed for conditional ranking protocols (e.g., DPO), our proposed preference acquisition protocol introduces Dove, a new preference optimization objective that upweights the joint probability of the chosen instruction-response pair over the rejected instruction-response pair. Interestingly, we find that the LLM trained with joint instruction-response preference data using Dove outperforms the LLM trained with DPO by $5.2\%$ and $3.3\%$ win-rate for the summarization and open-ended dialogue datasets, respectively. Our findings reveal that joint preferences over instruction and response pairs can significantly enhance the alignment of LLMs by tapping into a broader spectrum of human preference elicitation. The data and code is available at https://github.com/Hritikbansal/dove.
1 Introduction
Recently, alignment [42, 31] has emerged as a crucial step in enhancing the performance of large language models (LLMs) [1, 30, 44, 3, 11, 48, 19] in diverse real-world applications [22, 65, 57, 12, 21]. In particular, the aligned LLMs can produce high-quality outputs that satisfy human values and intents, including helpfulness, coherence, and harmlessness [4, 31]. LLM alignment hinges on the quality of preferences acquired from human or AI annotators [31]. Among the various preference acquisition protocols [23, 58, 39, 8], the ranking-based approach is the most widely used paradigm for aligning LLMs[42, 31, 6, 49, 45]. Specifically, in this approach the annotator has to compare a pair of responses conditioned on a fixed context. For instance, human can select a ‘preferred’ response by comparing a pair of responses for the instruction ‘Create a list of four fruits other than Apple’ (Figure 1 (left)).
Besides ranking preferences conditioned on a fixed context, humans can also express preferences in non-identical contexts. For example, while browsing reviews for products on an e-commerce website, humans are likely to prefer an accurate and detail-oriented review for a camera over an incoherent, vague movie review even though the products (camera and movie) are qualitatively different. Although the traditional conditional rankings provide rich preference for alignment, they fail to holistically capture the various dimensions of reasoning of human preferences. In this work, we revisit the traditional paradigm of conditional preference acquisition and propose a new approach for jointly eliciting preferences over instruction-response pairs. This method aims to uncover diverse reasoning paths in the process of acquiring feedback.
In this work, we develop a framework to acquire preferences jointly over instruction-response pairs. Starting from an instruction-response data consisting of response $R_{i}$ for instruction $I_{i}$ (say $i\in\{1,2\}$), we acquire ranking-based preferences over the instruction-response pairs $(I_{1},R_{1})$ and $(I_{2},R_{2})$. As shown in Figure 1 (right), we aim to understand whether the response in the pair $X$ is perceived better than the response in the pair $Y$. For instance, humans would prefer a helpful response to the instruction ‘Create a list of four fruits’ over a response that completely ignores the instruction ‘Create a list of beach activities’. This suggests that we can reveal preference axes like adherence to instructions, grammatical fluency, and clarity, even when following a joint preference optimization protocol. In addition, our protocol can elicit human preference behaviours that are obfuscated in the prior protocols, and redefines conditional preference elicitation as a special case where the instructions are identical.
Figure 1: Overview of the proposed Joint Preference Optimization paradigm. (Left) We show that the conditional preference acquisition method would require the annotators to compare two responses for an identical instruction. (Right) We show that the annotators can also assign rankings jointly over instruction-response pairs. Specifically, the annotator prefers a helpful response (e.g., Apple … Grape) over a response that ignores the context of the instruction (e.g., wear sunscreen … litter). Our framework thus elicits preferences that are obfuscated in the prior approach.
Next, we propose Dove, a framework for aligning LLMs with our proposed joint preference elicitation scheme. Specifically, it upweights the joint probability of the chosen instruction-response pair over the rejected instruction-response pair. This differs from the other frameworks that assume conditional rankings in their feedback data, such as DPO [35, 5] and preference optimizations that train a separate reward model such as PPO and rejection sampling [40, 29]. Further, we further point that Dove subsumes the prior preference optimizations as conditional rankings are a special case of joint preferences (e.g., when $I_{1}=I_{2}$). In our experiments, we focus on extending and comparing against DPO because of their simplicity, stability, and high-performance. However, our framework can be easily applied to PPO-like approaches by training a reward model on the joint preferences.
Finally, we conduct experiments to explore the new reasoning paths enabled by joint preference elicitation, followed by aligning LLMs with the Dove objective. To do so, we explore the interplay between the feedback data collected under conditional rankings and joint preferences protocol. In addition, we ask human annotators to explain their preference decisions, uncovering new reasoning paths that highlight the complexities of the preference acquisition process (§4). After feedback acquisition, we aim to investigate the impact of diverse preferences collected from conditional and joint preferences on LLM alignment. In our experiments, we align a Mistral-7B LLM with the preferences acquired from the conditional rankings and joint preferences, using our Dove algorithm. We find that the Dove outperforms the supervised finetuned LLM by $30\%$ and $18\%$ win-rate against the gold responses on the unseen instructions from the summarization and open-ended dialogues datasets, respectively. Surprisingly, we find that Dove can effectively tap into the diverse preferences in the conditional and joint feedback data and outperforms DPO by $5.2\%$ and $3.3\%$ win-rate points on the summarization and open-ended dialogues, respectively. This indicates that by utilizing the diverse preference signals present in the existing data, we can align an LLM robustly without acquiring additional instruction-response data.
2 Background
In this work, our aim is to align language models to generate outputs that are preferred by humans across various dimensions such as helpfulness and coherence. The process of aligning a base model, which is pretrained on a large corpus of text [13, 38, 41, 33], involves multiple steps: (a) instruction-response data collection, (b) supervised fine-tuning, (c) preference data acquisition, and (d) deployment of an alignment algorithm. The instruction-response data can be either hand-crafted by humans [14, 54] or generated by machines [43, 49]. Subsequently, the base model undergoes supervised fine-tuning (SFT) on the instruction-response pairs [66, 56, 54, 34, 59, 17, 60, 55, 62, 47]. Following SFT, feedback data is acquired under a specific acquisition protocol (e.g., rankings) from the annotators (§2.1). Finally, an alignment algorithm trains the SFT model on the feedback data (§2.2).
2.1 Ranking Feedback Acquisition Protocol
Assume a supervised finetuned language model $p_{\text{sft}}$ that is capable of responding to user instructions (e.g., imperative tasks or questions). The goal of alignment is to ensure that the SFT model generates high-quality outputs, preferred by humans. To do so, we consider a set of instructions =$\mathcal{I}=\{I_{1},\ldots,I_{n}\}$ where $n$ is the number of instructions. Further, we consider a set of responses $\{R_{j}^{1},R_{j}^{2},\ldots,R_{j}^{k}\}$ where $k$ is the number of responses for each of the instruction $I_{j}\in\mathcal{I}$. This forms a dataset of instructions and their corresponding responses, $\mathcal{D}=\{(I_{j},{R_{j}^{1},R_{j}^{2},\ldots,R_{j}^{k}})\}$. [^1] We will drop the iterator over j when defining the dataset for the ease of notation. Next, we acquire conditional ranking-based feedback over the collected instruction-response data.
Under this feedback acquisition protocol, the annotator selects a chosen and rejected response from $\{R_{j}^{x},R_{j}^{y}\}$ conditioned on the instruction $I_{j}$ where $x,y\in\{1,2,\ldots,k\}$. The preference decision by the annotator is based on the perceived quality of the responses along various dimensions such as helpfulness (accuracy), coherence (grammar), and harmlessness (safety).
Formally, the annotator assigns a instruction-conditioned ranking feedback $c(I_{j},R_{j}^{x},R_{j}^{y})\in\{R_{j}^{x},R_{j}^{y},\text{Equal}\}$ where ‘Equal’ indicates that both responses are perceived equally good or bad. If $c(I_{j},R_{j}^{x},R_{j}^{y})=R_{j}^{x}$, this implies that the response $R_{j}^{x}$ is the chosen response while the $R_{j}^{y}$ is the rejected response by the annotator. As a result, the ranking protocol creates a conditional pairwise feedback data $\mathcal{D}_{C}=\{(I_{j},R_{j}^{x},R_{j}^{y},c(I_{j},R_{j}^{x},R_{j}^{y}))\}$. Next, we apply an alignment algorithm on this data to elicit human-preferred responses from the LLM.
2.2 Alignment Algorithm
Rafailov et al. [36] introduced direct preference optimization (DPO) that can align a language model without utilizing on an external reward model. Specifically, DPO requires that feedback data should consist of conditional preferences between a pair of responses for a given instruction. Additionally, the algorithm assumes a preference dataset $\mathcal{D}_{C}$ and the reference model $p_{\text{ref}}$ which is usually the supervised finetuned language model $p_{\text{sft}}$. Specifically, it aims to train an aligned model $p_{\theta}$ using an optimization objective that upweights the conditional probability of the chosen response $p_{\theta}(R_{j}^{w}|I_{j})$ over the rejected response $p_{\theta}(R_{j}^{\ell}|I_{j})$ where $R_{j}^{w}$ and $R_{j}^{\ell}$ are the chosen and rejected response, respectively. Formally, the optimization objective for DPO minimizes:
$\mathcal{L}_{\textsc{DPO}}(\theta;\mathcal{D}_{C},\beta,p_{\text{ref}})={\mathbb{E}}_{(I_{j},R_{j}^{w},R_{j}^{\ell})\sim\mathcal{D}_{C}}\left[\operatorname{log}\left(\sigma\left(\beta\operatorname{log}\frac{p_{\theta}(R_{j}^{w}|I_{j})}{p_{\text{ref}}(R_{j}^{w}|I_{j})}-\beta\operatorname{log}\frac{p_{\theta}(R_{j}^{\ell}|I_{j})}{p_{\text{ref}}(R_{j}^{\ell}|I_{j})}\right)\right)\right]$ (1)
where $\sigma$ denotes the sigmoid function and $\beta$ is a hyperparameter. Post-alignment, the model generates high-quality outputs for unseen instructions that are preferred by the annotators.
3 Joint Preference Optimization using Dove
3.1 Joint Preference Acquisition Protocol
In §2.1, we describe a common technique for feedback data acquisition that requires the annotators to assign a preferred and non-preferred label to a pair of responses for an instruction. However, this paradigm does not capture the complex and multidimensional aspects of human preferences [20, 46]. Specifically, the reasoning paths for making preference decisions depend upon the context in which the comparison is made. While the traditional ranking protocol compares the two responses under a fixed context, humans can perform pairwise comparisons jointly over instruction-response pairs. For example, consider two summaries, A and B, for articles X and Y, respectively; then, a human can reason and choose the response that better summarizes its corresponding article. Hence, it is critical to align language models with diverse feedback signals to elicit high-quality responses that humans prefer under various contexts.
In our setup, the annotator has to decide a chosen and rejected instruction-response pair $(I_{a},R_{a},I_{b},R_{b})$ where $R_{a}$ and $R_{b}$ are responses to the instructions $I_{a}$ and $I_{b}$, respectively, and $(I_{a},R_{a}),(I_{b},R_{b})\in\mathcal{D}$. We note that our joint preference setup is equivalent to the original ranking protocol when $I_{a}=I_{b}$. As before, the preference reasoning from the annotator will be based on subjective dimensions like helpfulness, coherence, and harmlessness. Formally, the annotator assigns a joint ranking feedback $h(I_{a},R_{a},I_{b},R_{b})\in\{(I_{a},R_{m}),(I_{b},R_{b}),\text{Equal}\}$ where ‘Equal’ indicates that both the instruction-response pairs are perceived equally good or bad. Finally, the joint preference optimization creates a pairwise feedback data $\mathcal{D}_{H}=\{(I_{a},R_{a},I_{b},R_{b},h(I_{a},R_{a},I_{b},R_{b}))\}$.
Our formulation suggests that we can obtain large-scale and diverse preference data (covering all possible combinations of $(I_{a},R_{a})$ and $(I_{b},R_{b})$) without the need for gathering additional instruction and response data, which is typically more difficult and costly to acquire. In addition, joint preference acquisition does not necessitate the presence of multiple responses for a given instruction that can be hard to collect for low-resource languages (e.g., Kalamang  [^2] https://endangeredlanguages.com/lang/1891?hl=en). Specifically, one can collect an instruction-response data $\mathcal{D^{\prime}}=\{(I_{a},R_{a})\}_{a=1}^{a=n}$, and acquire preferences on various combinations of instruction-response pairs. Finally, we assess the interplay between the joint feedback dataset $\mathcal{D}_{H}$ with the conditional feedback dataset $\mathcal{D}_{C}$ along with qualitative examples in §4.
3.2 Dove
Here, we propose Dove, a preference optimization objective that learns to align the language models with the preferences acquired jointly over the instruction-response pairs. We assume a joint preference dataset $\mathcal{D}_{X}=\{(I_{i}^{w},R_{i}^{w},I_{j}^{\ell},R_{j}^{\ell})\}$, that can be constructed from $\mathcal{D}_{H}$, where $(I_{i}^{w},R_{i}^{w})$ and $(I_{j}^{\ell},R_{j}^{\ell})$ are the chosen and rejected instruction-response pairs, respectively. Similar to DPO, we start with a reference model $p_{\text{ref}}$ which is usually the supervised finetuned language model $p_{\text{sft}}$. Specifically, the Dove objective aims to learn an aligned model $p_{\theta}$ by upweighting the joint probability of preferred responses $p(R_{i}^{w},I_{i}^{w})$ over non-preferred responses $p(R_{j}^{\ell},I_{j}^{\ell})$. Formally, the optimization objective for Dove minimizes:
$\mathcal{L}(\theta;\mathcal{D}_{X},\beta,p_{\text{ref}})$ $=\mathbb{E}_{(I_{j}^{w},R_{j}^{w},I_{j}^{\ell},R_{j}^{\ell})\sim\mathcal{D}_{X}}\left[\log\left(\sigma\left(\beta\log\frac{p_{\theta}(R_{i}^{w},I_{i}^{w})}{p_{\text{ref}}(R_{i}^{w},I_{i}^{w})}-\beta\log\frac{p_{\theta}(R_{j}^{\ell},I_{j}^{\ell})}{p_{\text{ref}}(R_{j}^{\ell},I_{j}^{\ell})}\right)\right)\right]$ (2)
where $\sigma$ denotes the sigmoid function and $\beta$ is a hyperparameter. Further, we show that Eq. 3 reduces to the DPO formulation (Eq. 1) when the instructions $I_{i}=I_{j}$ in Appendix §D. We can also see that the Dove objective aims to learn an aligned model $p_{\theta}$ by upweighting the conditional probability of preferred responses $p(R_{i}^{w}|I_{i}^{w})$ over non-preferred responses $p(R_{j}^{\ell}|I_{j}^{\ell})$, along with a correction factor based on the prior probability of the instructions under the language model $p_{\theta}(I_{i}^{w})$ and $p_{\theta}(I_{j}^{\ell})$. In §5, we utilize Dove to align language models to generate human-preferred summaries and answer open-ended instructions.
4 Interplay between Feedback Protocols
4.1 Instruction-Response Acquisition
The instruction-response data is a collection of real-world queries that are presented to the text AI assistants. In this work, we consider two kinds of instruction-response data. First, we consider a filtered version of the TL;DR summarization dataset [51] from [42] consisting of Reddit posts, their summarizes, and human preferences over a pair of summaries for a given post. Throughout the dataset, the task is of summarization that is close-ended and well-defined for language models. Second, we consider the single-turn dialogues from the helpful-base subset of the Anthropic-HH dataset [7]. Specifically, this dataset consists of open-ended instructions with a collection of responses ranging from ‘Which coffee bean is better for a morning roast?’ to ‘How do I attract more hummingbirds in my yard?’.
Both these datasets come with the train and test split where each instance consists of an instruction and a pair of responses $\mathcal{D}=\{(I_{i},R_{i}^{1},R_{i}^{2})\}_{i=1}^{n}$ where $n$ is the dataset size. In this work, we collect AI and human feedback on the instruction-response data from their train split and filter instances where instructions are repeated. We can directly compare the two responses for the fixed instruction and construct a ranking feedback dataset $\mathcal{D}_{C}=\{(I_{i},R_{i}^{1},R_{i}^{2},c(I_{i},R_{i}^{1},R_{i}^{2}))\}$. To acquire preferences jointly over the instruction-response pairs, we select one of the responses, at random, from every instance of $\mathcal{D}$ to construct $\mathcal{D}_{S}=\{(I_{i},R_{i})\}$ where $R_{i}\in\{R_{i}^{1},R_{i}^{2}\}$. Subsequently, we create the joint instruction-response pairs by matching every instance $(I_{i},R_{i})\in\mathcal{D}_{S}$ with another instance $(I_{j},R_{j})\in\mathcal{D}_{S}$ to get $\mathcal{D}_{H}=\{(I_{i},R_{i},I_{j},R_{j},h(I_{i},R_{i},I_{j},R_{j}))\}$ of the same size as $\mathcal{D}_{S}$ and $\mathcal{D}_{C}$. In §5, we will utilize $\mathcal{D}_{S}$ to SFT the base model, and $\mathcal{D}_{C}$ and $\mathcal{D}_{H}$ as preference datasets for LLM alignment. We provide the dataset statistics in Appendix §E.
4.2 Feedback from AI and Humans
Feedback from AI.
Prior work [15, 7] has shown that AI feedback can be leveraged to align language models to generate helpful and harmless responses to unseen instructions. In addition, acquiring AI feedback at large-scale is more accessible and cheaper in comparison to human feedback. To this end, we collect feedback over a pair of responses for a fixed instruction, and joint instruction-response pairs without identical instructions from GPT-3.5-Turbo-0125 (ChatGPT). The choice of ChatGPT was motivated by its affordability (e.g., output tokens from ChatGPT are 50×50$\times$50 × cheaper than GPT-4).
To collect ranking feedback over a pair of responses for a fixed instruction, we prompt ChatGPT to assign a chosen response. To mitigate any bias from the ordering of the two responses, we run two queries for all comparisons to cover all possible orderings. When the ChatGPT preferences flip by flipping the order of the two responses, then we consider it a tie, similar to [8, 9]. Specifically, the AI is instructed to provide its preference based on the accuracy, coherence, and harmlessness of the responses.
To collect AI preferences jointly over the instruction-response pairs, we prompt ChatGPT to decide the response that better answers its corresponding instruction. Similar to the previous scenario, we run two queries for all comparisons to mitigate any ordering bias and provide guidelines to choose the response that is more accurate, coherent, and harmless. We collected approximately $50K$ comparisons across both feedback acquisition protocols for the summarization and Anthropic-Helpful dataset, at a cost of \$$\$100$. We provide the AI prompts in Appendix §F.
Feedback from Humans.
In this work, we also collect human preferences for $2000$ comparisons over summarization and Anthropic-Helpful dataset. Such a data is useful for providing insights into the human behavior under different preference acquisition protocols (§4.3). In addition, this data aids in agreement between the ChatGPT and human decisions.
Specifically, we ask two annotators to assign a chosen response or choose ‘equal’ after comparing the quality of the responses along the same dimensions as ChatGPT guidelines. The human annotations were collected from Amazon Mechanical Turk (AMT) from the participants that passed a preliminary qualification exam. In total, we spent \$$\$720$ on human feedback acquisition. We provide the screenshot of the annotation UI in Appendix §G. We find that the AI achieves an agreement score of $64\%$ with the human annotators in our agreement analysis, compared to the $69\%$ inter-annotator agreement (Section C). This highlights the high quality of the AI feedback across datasets and preference protocols (i.e., conditional and joint preferences over non-identical instructions).
4.3 Interplay Analysis
Setup.
Here, we aim study the interaction between the conditional rankings and joint rankings over non-identical instructions. Formally, each instruction-response pair $(I_{i},R_{i}^{x})$ from the conditional pairwise feedback dataset $\mathcal{D}_{C}$ where $x\in\{1,2\}$ can be assigned a preference $\mathcal{P}_{C}(I_{i},R_{i}^{x})$ among {‘chosen’, ‘reject’, ‘equal’}. For instance, $\mathcal{P}_{C}(I_{i},R_{i}^{1})=`\text{chosen}^{\prime}$ and $\mathcal{P}_{C}(I_{i},R_{i}^{2})=`\text{reject}$ if the response $R_{i}^{2}$ is rejected in the dataset $\mathcal{D}_{C}$ i.e., $c(I_{i},R_{i}^{1},R_{i}^{2})=R_{i}^{1}$. Similarly, we can assign a preference $\mathcal{P}_{H}(I_{i},R_{i})$ among {‘chosen’, ‘reject’, ‘equal’} to an instruction-response pair $(I_{i},R_{i})$ from the joint preference dataset $\mathcal{D}_{H}$. For instance, $\mathcal{P}_{H}(I_{i},R_{i})=`\text{chosen}^{\prime}$ and $\mathcal{P}_{H}(I_{j},R_{j})$ where $i!=j$ if the instruction-response pair $(I_{i},R_{i})$ is chosen in the dataset $\mathcal{D}_{H}$ i.e., $h(I_{i},R_{i},I_{j},R_{j})=R_{i}$.
To study the interplay between the preference protocols, we assess $\mathcal{P}_{C}(I_{i},R_{i})$, $\mathcal{P}_{C}(I_{j},R_{j})$, $\mathcal{P}_{H}(I_{i},R_{i})$ and $\mathcal{P}_{H}(I_{j},R_{j})$ for all $(I_{i},R_{i},I_{j},R_{j})\in\mathcal{D}_{H}$. Here, if $\mathcal{P}_{H}(I_{i},R_{i})=`\text{chosen}^{\prime}$ then $\mathcal{P}_{H}(I_{i},R_{i})=`\text{reject}^{\prime}$. For instance, if $\mathcal{P}_{C}(I_{i},R_{i})=`\text{chosen}^{\prime}$ and $\mathcal{P}_{C}(I_{j},R_{j})=`\text{chosen}^{\prime}$ then it implies that the annotators can reason about the joint preferences over a pair of instruction-response pairs that are originally preferred under the conditional ranking feedback protocol. We quantify study the interplay between the two ranking-based feedback from AI and Human annotators over summarization and open-ended Anthropic-Helpful datasets.
Results.
Data (Annotator) Decisive Indecisive
TL;DR (AI) 63.7% 36.2%
TL;DR (Human) 73.8% 25.7%
Anthropic-Helpful (AI) 68.5% 31.5%
Anthropic-Helpful (Human) 77.9% 22.0%
Average 71.0% 29.0%
(a) Results for the preferences acquired jointly over the instruction-response pairs where both the responses were either chosen or rejected under the conditional rankings protocol. Here, decisive implies that the annotators could assign a preference to one instruction-response pair over the other.
Data (Annotator) C $>$ R C $<$ R Indecisive
TL;DR (AI) 53.3% 14.3% 30.4%
TL;DR (Human) 41.6% 22.2% 36.1%
Anthropic-Helpful (AI) 54.5% 17.6% 27.8%
Anthropic-Helpful (Human) 57.1% 21.4% 21.4%
Average 52.0% 19.0% 29.0%
(b) Results for the preferences acquired jointly over the instruction-response pairs where one of the instruction-response pair was chosen (C) and the other pair was rejected (R) under the conditional rankings. Here, $C<R$ implies that the instruction-response pair that was rejected under conditional rankings is actually preferred over an instruction-response pair that was rejected under the conditional rankings.
Table 1: Results for the interplay between the conditional rankings and joint rankings acquired over non-identical instructions for summarization and open-ended dialogue dataset. In total, we compare $48K$ and $1K$ annotations from the AI and humans, respectively.
We present the results for the interaction analysis in Table 1. In Table 1(a), we study the joint preferences over the instruction-response pairs $(I_{i},R_{i},I_{j},R_{j})$ where the individual instruction and response data is either chosen or rejected in the conditional feedback protocol (e.g., $\mathcal{P}_{C}(I_{z},R_{z})=`\text{chosen}^{\prime}$ for $z\in\{i,j\}$). Interestingly, we find that the annotators can assign a decisive preference (e.g., $(I_{i},R_{i})>(I_{j},R_{j})$) in $71\%$ of the joint comparisons. While we observe that the annotators assign a ‘tie’ to $29\%$ of the comparisons. This highlights the existence of valid preference decisions that remained obfuscated in the traditional approach for ranking-based feedback acquisition.
In Table 1(b), we study the joint preference over the instruction-response pairs $(I_{i},R_{i},I_{j},R_{j})$ where one of them is chosen and the other is rejected in the conditional feedback protocol (e.g., $\mathcal{P}_{C}(I_{i},R_{i})=`\text{chosen}^{\prime}$ and $\mathcal{P}_{C}(I_{i},R_{i})=`\text{reject}^{\prime}$). To our surprise, we find that the annotators do not prefer the instruction-response pair that was chosen under the conditional feedback protocol in $48\%$ of the comparisons. Specifically, there are $19\%$ of the comparisons where rejected pair (R) is preferred over the chosen pair (C) and $28\%$ of the comparisons where the annotators considered the pair equally good or bad. This highlights that both human and AI annotators’ perceptions of preferred and non-preferred data depends on the context of the comparisons, indicating that feedback acquisition is a multifaceted phenomenon.
Qualitative Examples.
To probe the reasoning paths of the human annotators used for decision making, we ask them to provide brief explanations for their feedback decisions regarding a few conditional and joint preferences. We provide a list of qualitative examples consisting of instructions, responses, and respective preferences in Appendix §H. In Figure 8, we discovered that human annotators provided decisive feedback when comparing instruction-response pairs, basing their decisions on the accuracy of the responses. In Figure 11, we find that the human annotators preferred a instruction-summary pair, that was rejected under the conditional preference, because it provides a fuller picture of the original reddit post. In summary, we expose the multi-faceted reasoning paths of humans in joint instruction-response feedback acquisition that would have been concealed in the conditional feedback acquisition paradigm.
5 LLM Alignment
In the previous sections, we show that the humans and AI are capable of providing ranking-based feedback for a pair of responses for identical and non-identical instructions in the context. Here, we aim to study the impact of feedback data on eliciting high-quality responses from the large language models.
5.1 Setup
Here, we aim to align Mistral-7B  [19], a strong base LLM for its model capacity. We experiment with two datasets that exhibit diverse characteristics: (a) TL;DR dataset where the instruction is to summarize Reddit posts, and (b) open-ended dialogues from Anthropic-Helpful dataset (§4.1). In particular, we collect a conditional preference data $\mathcal{D}_{C}$ and joint preference data for non-identical instructions $\mathcal{D}_{H}$ of similar data sizes on both the datasets from ChatGPT. Then, we convert the conditional preference data into an instruction-response data for supervised finetuning $\mathcal{D}_{\text{SFT}}$.
First, we supervise finetune the entire base LLM model parameters with the SFT dataset to ensure that the preference data is in-policy for the alignment algorithms  [36]. Subsequently, we apply DPO algorithm on the SFT model using the conditional preference data for $10$ epochs and $5$ epochs for the summarization and Anthropic-helpful data, respectively. Specifically, we use low-rank adaptation  [18] of SFT model during DPO alignment. The DPO optimization was trained on a single GPU Nvidia A6000 with a batch size of $32$.
We note that our proposed Dove algorithm can utilize both the conditional preferences and joint preference with non-identical context. It is because the conditional preferences can be viewed as joint preferences with identical context. As a result, we train the base LLM with Dove algorithm after merging conditional and joint preferences data $\mathcal{D}_{M}=\mathcal{D}_{C}\cup\mathcal{D}_{H}$. We keep the hyperparameters (e.g., $\beta$), number of epochs, and the batch size identical to the DPO algorithm. In our experiments, we also train Dove algorithm on the joint preferences with non-identical instructions and highlight their usefulness for LLM alignment. We provide more details on training setup in Appendix §I.
Post-alignment, we evaluate the aligned model responses against the gold responses in the dataset’s test split. Specifically, both datasets come with a human-preferred response for an instruction, which is treated as the gold response. We utilize ChatGPT to compare model and gold responses to decide on the preferred response or a tie. Finally, we report the win-rate of the model responses as the evaluation metric for $500$ unseen instructions  [36].
Figure 2: Results for aligning LLMs with the Dove preference optimization objective. We compare the win-rate against the gold responses of the supervised finetuned (SFT), DPO-aligned and Dove-aligned LLM on the (a) TL;DR summarization and (b) the Anthropic-Helpful datasets. In our experiments, we utilize ChatGPT to compare the model responses with the gold responses. We generate model responses for three sampling temperatures. The results are averaged over three runs of the preference optimization objectives.
5.2 Results
We compare the performance of the SFT, DPO, and Dove aligned models in Figure 2. In particular, we report the win-rate against the gold responses for the model responses generated from various sampling temperatures $\text{T}\in\{0.001,0.5,1.0\}$.
Dove outperforms SFT model.
We find that the Dove achieves high win-rates across all sampling temperatures. Specifically, we observe that Dove outperforms the SFT model by $29.1\%$ and $18\%$ on the close-ended summarization and open-ended dialogue dataset, respectively, averaged across the sampling temperatures. This indicates that Dove can utilize the diverse set of feedback from the conditional and joint preferences to align LLMs.
Dove outperforms DPO algorithm.
Further, we aim to understand whether Dove is able to tease out useful feedback signals from the combination of the conditional preferences and joint preferences over instruction-response pairs. Surprisingly, we find that Dove outperforms DPO by $5.2\%$ and $3.3\%$ win-rate points on the summarization and helpfulness datasets, respectively. In addition, the performance of Dove is better than Dove across all the sampling temperatures. This highlights that one can improve the alignment of the LLMs by leveraging novel preference acquisition paths without collecting new instruction-response data. Further, our results indicate that Dove is a robust alignment algorithm that can elicit high-quality outputs by learning from diverse ranking-based preferences.
Figure 3: Win-rate against the gold response in the TL;DR and Anthropic-Helpful datasets averaged over three sampling temperatures. We study the impact of the joint preferences over non-identical instructions using Dove.
Impact of Joint Preferences over Non-Identical Instructions.
Here, we aim to understand the sole impact of joint preferences acquired over non-identical instructions on the performance of the Dove algorithm. To do so, we train Dove algorithm with joint feedback data $\mathcal{D}_{H}$ only. We present the results averaged across the three sampling temperatures in Figure 3. We find that training with joint preferences over non-identical instructions achieves $71.7\%$ and $69\%$ win-rate on the summarization and anthropic-helpful datasets, respectively. This indicates that it is possible to align LLMs with just joint preferences over instruction-response data without any conditional preferences too. Furthermore, this highlights that the feedback paths exposed in our setup are robust and effective for eliciting preferred model responses. We also show that the training with joint preferences scales with the amount of feedback data using the Dove algorithm in Appendix §J.
6 Related Work
Alignment using Reinforcement Learning.
Large Language Models (LLMs) showcase impressive few-shot performance [11] which is further improved by instruction tuning followed by alignment with preferences data [50, 31, 15, 66, 40, 53]. This alignment is usually done by first optimizing for a reward model on a preference data [10], followed by aligning the LLMs distribution to maximize the learned reward model (using RLHF/RLAIF). In this work, we show that preference acquisition is a complex phenomenon and elicits joint preferences over instruction-response data. Further, we study the interactions between joint preferences and the traditional approach of conditional rankings for preference elicitation.
Reward Free Policy Alignment.
[37] introduced Direct Preference Optimization (DPO) that mitigates the instability due to PPO for reward maximization by skipping reward modeling and optimizing directly within the model parameter space. [32] adds an explicit penalty term to avoid a reduction in the likelihood of preferred pairs. [16] improves over DPO by using human-aware loss based on prospect theory. Concurrent to our work, [61] explores the comparison between two responses for non-identical instructions. However, they do not uncover new preference acquisition paths, and instead rely on augmenting the existing conditional rankings. In addition, we introduce Dove, a novel preference optimization, complementary to existing works, that learns from joint preferences over instruction-response data and achieves good performance on diverse tasks. We refer to §B for a detailed related work.
7 Conclusion
In this work, we propose a framework that elicits preferences jointly over instruction-response pairs. Further, we find that the joint preference optimization uncovers new paths of human reasoning that remain obscured in the traditional approach. Additionally, we propose Dove, a novel preference optimization objective for aligning LLMs. In our experiments, we show that it outperforms DPO on summarization and dialogue datasets. We note that the number of joint preferences over instruction-response data scales quadratically with the number of instances in the instruction-response dataset. Therefore, identifying the most informative joint comparisons for robust LLM alignment represents a relevant area for future research. While traditional LLM evaluation has focused on conditional rankings, LLM evaluation through joint rankings would be an important future work.
Appendix A Limitations
While there are various protocols for feedback acquisition, our work is focused on acquiring rankings on a pair of responses under a fixed context or jointly over instruction-response pairs. While ranking-based protocol is widely accepted, there are several limitations associated with it. For instance, conditional or joint rankings do not quantify the strengths or weaknesses for a particular task. In addition, [8] show that different forms of feedback data often disagree with each other. This highlights at the complex and multidimensional aspects of human preferences.
In our work, we propose the joint acquisition of feedback for pairs of instruction-response over diverse tasks (e.g., comparing a movie review with an e-commerce product review). However, acquiring joint preferences may be challenging for certain combinations of instruction-response data. This difficulty arises particularly when the distributions of the instructions are significantly dissimilar. For example, it may be challenging to compare feedback for a response to the instruction ’how to cook fried rice?’ with a response to ’how to steal my neighbor’s wifi?’. In this scenario, the first instruction aims to elicit a helpful response, while the latter seeks a harmful one. In such cases, it is reasonable to expect that human annotators will be biased, preferring more helpful responses over harmful ones or vice versa. Therefore, introducing a notion of instruction similarity to decide which instruction-response pairs to compare under the joint preference protocol might be beneficial.
Finally, we acquire human annotations from Amazon Mechanical Turk (AMT) where most of the annotators belong to the U.S. or Canada regions. Hence, the preferences in our dataset are not represented of the diverse demographics in the world. It is pertinent that the future work should study the impact of the diverse groups on the feedback data behaviours and subsequent LLM alignment [63].
Appendix B Detailed Related Work
Alignment using Reinforcement Learning.
Large Language Models (LLMs) showcase impressive zero-shot and few-shot performance [11] with instruction tuning further improving their downstream performance [50]. Not all responses for a prompt are preferable due to a variety of reasons, thus aligning them with human preference is now a common step in fine-tuning the pipeline. This alignment is usually done by first optimizing for a reward model on preference data [10, 24, 8], followed by aligning the LLMs distribution that maximizes the learned reward model using Reinforcement Learning (RLHF) [40, 31], with optional Divergence penalty [53] to avoid deviating from the reference policy. Despite the success of RLHF, collecting human feedback can be challenging, however, [15, 28, 66] observe that LLMs can provide feedback motivating Reinforcement Learning through AI feedback (RLAIF). In this work, we show that preference acquisition is a complex phenomenon and elicit joint preferences over instruction-response data. Further, we study the interactions between joint preferences and traditional approach of conditional rankings for preference elicitation.
Reward Free Policy Alignment.
[37] introduced Direct Preference Optimization (DPO) that mitigates the instability due to PPO for reward maximization by optimizing directly within the model parameter space, hence by-passing the reward modeling step. [26] extends this framework where instead of two responses, alignment is done over the list of responses while [25] improves DPO using statistical rejection sampling. [2] provides an offset in the DPO objective to increase the margins and [32] suggests adding an explicit penalty term to avoid a reduction in the likelihood of preferred pairs over the DPO training. Contrary to our work where compare the joint distributions, [61] proposes RPO that compares the conditional likelihood of a winning response with the losing response of another prompt. Beyond DPO, more recently [16] proposed a human-aware loss function-based framework using prospect theory named KTO, and [5] proposes IPO that uses human preferences expressed as pairwise preferences. Lastly, [64] uses sequence likelihood calibration to align the model from human preference. Here, we introduce Dove, a novel preference optimization, complementary to existing works, that learns from joint preferences over instruction-response data and achieves good performance on diverse tasks.
Appendix C Agreement Analysis.
Here, we compare the agreement between the human and AI preference decisions for the summarization and open-ended tasks under the two acquisition protocols. To establish the gold chosen response in the human annotations, we compare the preference from the two annotators. If the two annotators agree in their response (e.g., Response $R_{1}$ $>$ Response $R_{2}$), then it is considered the gold chosen response. If the annotators disagree in their preferences (Response $R_{1}$ $>$ Response $R_{2}$ for annotator 1 and A$<$B for annotator 2) then we consider it as tie. If an annotator is indecisive while the other annotator makes a decisive preference (e.g., Response $R_{1}$ $=$ Response $R_{2}$ for annotator 1 and A$>$B for annotator 2), then the gold response is the one for which a decisive feedback has been provided i.e., Response $R_{1}$ In addition, the inter-annotator agreement is computed by comparing the preference from the two annotators on the identical comparison. In case of disagreements where one of the annotator chooses ‘equal’, the agreement score is considered as $0.5$. We apply similar guidelines to compute human-AI agreement.
Dataset Ranking Protocol Human-Human Human-AI
Summarization (TL;DR) Two Responses under Fixed Instruction 69% 63%
Anthropic-Helpful 70.1% 62%
Summarization (TL;DR) Joint Instruction and Response Pairs 62% 60%
Anthropic-Helpful 74.8% 71%
Average 69% 64%
Table 2: Agreement analysis between within human annotators and gold human feedback and AI (ChatGPT) feedback. We perform the agreement calculations for the two ranking protocols: (a) conditional rankings, and (b) joint preferences over instruction-response pairs where instructions are non-identical. In addition, we assess the agreement rates over the two datasets: (a) summarization and (b) Anthropic-helpful dataset.
We present the agreement results in Table 2. We find that the average agreement $69\%$ and $64\%$ between the human-human and human-AI annotators, respectively. These agreement scores are close to the annotation agreements presented in prior works [22, 8]. Interestingly, the agreement scores vary based on the underlying distribution of the instruction-response pairs and the choice of ranking protocol. Overall, our results highlight that AI provides high-quality feedback across summarization and open-ended dataset for both the feedback protocols.
Appendix D Proof for Dove subsuming DPO
Here we highlight a result that reduces Dove into DPO when the prompts are the same.
Lemma 1.
Under the case where $\mathcal{D}_{X}=\{(I_{i},R_{i},I_{i},R_{j})\}$, that is, prompts are the same for preferred and not-preferred prompt generation pairs, $\mathcal{L}_{\textsc{DPO}}(\theta;\mathcal{D}_{C},\beta,p_{\text{ref}})=\mathcal{L}_{\textsc{Dove}}(\theta;\mathcal{D}_{X},\beta,p_{\text{ref}})$, where $\mathcal{D}_{C}=\{(I_{j},R_{j}^{w},R_{j}^{\ell})\}$.
Proof.
$\mathcal{L}_{\textsc{Dove}}(\theta;\mathcal{D}_{X},\beta,p_{\text{ref}})$ $=\mathbb{E}_{(I_{j}^{w},R_{j}^{w},I_{j}^{\ell},R_{j}^{\ell})\sim\mathcal{D}_{X}}\left[\log\left(\sigma\left(\beta\log\frac{p_{\theta}(R_{i}^{w},I_{i}^{w})}{p_{\text{ref}}(R_{i}^{w},I_{i}^{w})}-\beta\log\frac{p_{\theta}(R_{j}^{\ell},I_{j}^{\ell})}{p_{\text{ref}}(R_{j}^{\ell},I_{j}^{\ell})}\right)\right)\right]$ (3)
$=\mathbb{E}_{(I_{j}^{w}R_{j}^{w},I_{j}^{\ell},R_{j}^{\ell})\sim\mathcal{D}_{H}}\left[\log\left(\sigma\left(\beta\log\frac{p_{\theta}(R_{i}^{w}|I_{i}^{w})p_{\theta}(I_{i}^{w})}{p_{\text{ref}}(R_{i}^{w}|I_{i}^{w})p_{\text{ref}}(R_{i}^{w}|I_{i}^{w})}\right.\right.\right.$
$\quad\left.\left.\left.-\beta\log\frac{p_{\theta}(R_{j}^{\ell}|I_{j}^{\ell})p_{\theta}(I_{j}^{\ell})}{p_{\text{ref}}(R_{j}^{\ell},I_{j}^{\ell})p_{\text{ref}}(I_{j}^{\ell})}\right)\right)\right]$
$={\mathbb{E}}_{(I_{j},R_{j}^{w},R_{j}^{\ell})\sim\mathcal{D}_{C}}\left[\operatorname{log}\left(\sigma\left(\beta\operatorname{log}\frac{p_{\theta}(R_{j}^{w}|I_{j})}{p_{\text{ref}}(R_{j}^{w}|I_{j})}-\beta\operatorname{log}\frac{p_{\theta}(R_{j}^{\ell}|I_{j})}{p_{\text{ref}}(R_{j}^{\ell}|I_{j})}\right)\right)\right]$ (4)
$=\mathcal{L}_{\textsc{DPO}}(\theta;\mathcal{D}_{C},\beta,p_{\text{ref}})$ (5)
The proof follows from applying bayes rule and substituting $I_{j}^{w}=I_{j}^{\ell}=I_{j}$. ∎
Appendix E Dataset Statistics
We present the dataset statistics in Table 3. We report the number of instructions after filtering the instances with repeated instructions. Each instance in the dataset consists of an instruction, and a pair of responses. Originally, the number of AI-generated conditional and joint preferences equals the number of instructions data. Here, we report the number of instances for which we observe a decisive preference from ChatGPT i.e., after removing the ties.
OpenAI TL;DR Summarization Dataset Number
Number of instructions 11.8K
Number of AI generated conditional preferences 7.2K
Number of AI generated joint preferences 7.7K
Anthropic-Helpful Dataset
Number of instructions 12.8K
Number of AI generated conditional preferences 9.4K
Number of AI generated joint preferences 8.5K
Table 3: Statistics for the train split of the summarization and open-ended dialogue datasets.
Appendix F ChatGPT Prompts
We present the ChatGPT for acquiring conditional rankings feedback and joint preferences over instruction-response pairs in Table 4 and Table 5, respectively.
You are a helpful following assistant whose goal is to select the preferred output for a given instruction. Answer the question by printing only a single choice from [”Output (a)”, ”Output (b)”] (without quotes) corresponding to the correct answer with no other text.
#### Annotation Guideline
In this task, we will ask you to select the preferred output AI model’s responses to instructions.
You will read examples, which are composed of the following:
1. an Instruction we give to the AI system
2. Output (a), the first output from the AI system
3. Output (b), the second output from the AI system
Your task is to decide which response is better for each example. There are several dimensions that you can think along. Consider the following dimensions:
Accuracy: The output sentence should be factually consistent with the instruction and the input. Coherence: The output sentence should be easy to understand and free of grammatical errors when read on its own. Harmlessness: The output sentence should not contain derogatory, harmful, or toxic connotations.
You should answer using only Output (a) or Output (b) depending on which response is better.
### Instruction:
${instruction}$
### Output (a):
${output_{1}}$
### Output (b):
${output_{2}}$
## Preferred Output:
Figure 4: GPT-3.5-Turbo API prompt for comparisons on identical instructions
You are a helpful following assistant whose goal is to select the preferred output for a given instruction. Answer the question by printing only a single choice from [”Output (a)”, ”Output (b)”] (without quotes) corresponding to the correct answer with no other text.
#### Annotation Guideline
In this task, we will ask you to select the preferred output AI model’s responses to instructions.
You will read examples, which are composed of the following:
1. Instruction (a), the first input to the AI system and Output (a), the first output from the AI system
2. Instruction (b), the second input to the AI system and Output (b), the second output from the AI system
Your task is to decide which response is better for each example. There are several dimensions that you can think along. Consider the following dimensions:
Accuracy: The output sentence should be factually consistent with the instruction and the input. Coherence: The output sentence should be easy to understand and free of grammatical errors when read on its own. Harmlessness: The output sentence should not contain derogatory, harmful, or toxic connotations.
You should answer using only Output (a) or Output (b) depending on which response is better.
### Instruction (a):
${instruction_{1}}$
### Output (a):
${output_{1}}$
### Instruction (b):
${instruction_{2}}$
### Output (b):
${output_{2}}$
## Preferred Output:
Figure 5: GPT-3.5-Turbo API prompt for comparisons on non-identical instructions
Appendix G Human Annotation Platform
We present the screenshots for the human interface in the Figure 6 (conditional rankings) and Figure 7 (joint ranking preferences over instruction-response pairs).
Figure 6: Human annotation interface for Conditional Rankings
Figure 7: Human annotation interface for joint preferences over instruction-response pairs.
Appendix H Qualitative Examples
In this section, we present the qualitative examples to study the interplay between the conditional rankings and the joint preference over instruction-response pairs. Here, we acquire ranking feedback from the human annotators and ask them to provide the reasoning for their decision.
H.1 Anthropic-Helpful Examples
We present the qualitative examples for the preferences acquired for the Anthropic-helpful dataset in Figure 8, 9, and 10. We present our observations in the figure captions.
Figure 8: Interplay between the conditional rankings and joint rankings and reasoning acquired from the human annotators for the Anthropic-Helpful dataset. In this example, we find that the response B and D are rejected under the conditional rankings. When asked to compare the response B and D, humans consider that the response B answers Instruction 1 better than response D answers Instruction 2. This indicates that the joint preference humans elicits a decisive feedback between two responses that were rejected under the conditional rankings.
Figure 9: Interplay between the conditional rankings and joint rankings and reasoning acquired from the human annotators for the Anthropic-Helpful dataset. In this example, we find that the response A and C are accepted under the conditional rankings. When asked to compare the response A and C, humans consider that the response A answers Instruction 1 better than response C answers Instruction 2. This indicates that the joint preference humans elicits a decisive feedback between two responses that were accepted under the conditional rankings.
Figure 10: Interplay between the conditional rankings and joint rankings and reasoning acquired from the human annotators for the Anthropic-Helpful dataset. In this example, we find that the response A is accepted and D is rejected under the conditional rankings. When asked to compare the response A and D, humans consider that the response A answers Instruction 1 better than response D answers Instruction 2. This indicates that a response that was preferred (rejected) under the conditional rankings can still be preferred (rejected) under the joint rankings.
H.2 TL;DR Summarization Examples
We present the qualitative examples for the preferences acquired for the TL;DR summarization dataset in Figure 11, 12, and 13. We present our observations in the figure captions.
Figure 11: Interplay between the conditional rankings and joint rankings and reasoning acquired from the human annotators for the TL;DR summarization dataset. In this example, we find that the response B is accepted and C is rejected under the conditional rankings. When asked to compare the response B and C, humans consider that the response C answers Instruction 2 better than response B answers Instruction 1. This indicates that a response that was preferred (rejected) under the conditional rankings can be rejected (preferred) under the joint rankings, further highlighting at the complex and multidimensional nature of human preferences.
Figure 12: Interplay between the conditional rankings and joint rankings and reasoning acquired from the human annotators for the TL;DR summarization dataset. In this example, we find that the response B and C are accepted under the conditional rankings. When asked to compare the response B and C, humans consider that the response B answers Instruction 1 better than response C answers Instruction 2. This indicates that the joint preference humans elicits a decisive feedback between two responses that were accepted under the conditional rankings.
Figure 13: Interplay between the conditional rankings and joint rankings and reasoning acquired from the human annotators for the TL;DR summarization dataset. In this example, we find that the response A is considered to be equally good as response B for the instruction 1. In addition, response C is rejected in comparison to the response D for the instruction 2. However, when asked to compare the response A and C, humans consider that the response C answers Instruction 2 better than response A answers Instruction 1. This highlights that a rejected response can be preferred over a non-rejected response under joint rankings.
Appendix I Alignment Training Details
I.1 Supervised Finetuning Details
We present the SFT details in  Table 4. We perform full-finetuning of Mistral-7B using the source code from https://github.com/abacaj/fine-tune-mistral.
Anthropic-Helpful Dataset
Learning Rate 1.5e-6
Batch Size 6
Epochs 3
OpenAI TL;DR Summarization Dataset
Learning Rate 2e-5
Batch Size 12
Epochs 3
Table 4: Training details for the supervised finetuning of Mistral-7B.
I.2 Dove
We present the training details for Dove preference optimization objective in the Table 5. We select the learning rate hyperparameter by sweeping over three learning rates: $\{1e-5,5e-5,5e-4\}$. We utilize the TRL library [52] for the DPO source code.
OpenAI TL;DR Summarization Dataset
Peak Learning Rate 5e-5
Optimizer AdamW [27]
Learning Schedule Cosine
Batch Size 32
Epochs 10
Warmup Steps 100
$\alpha$ (LoRA) 16
Dropout (LoRA) 0.05
Bottleneck $r$ (LoRA) 8
4bit Loading True
$\beta$ 0.1
Anthropic-Helpful Dataset
Peak Learning Rate 5e-5
Optimizer AdamW
Learning Schedule Cosine
Batch Size 32
Epochs 5
Warmup Steps 100
$\alpha$ (LoRA) 16
Dropout (LoRA) 0.05
Bottleneck $r$ (LoRA) 8
4bit Loading True
$\beta$ 0.1
Table 5: Training details for Dove preference optimization objective. We use the identical settings for DPO.
Appendix J Trends with Data Scaling
We aim to understand the impact of increasing the number of preferences collected jointly over instruction-response pairs, for non-identical instructions, on the win-rate against the reference summaries in the TL;DR summarization dataset using Dove algorithm. We present the results in Figure 14 for the sampling temperature of $0.001$. We find that the win-rate scales from $42.4\%$ to $71.7\%$ as the size of the dataset increases from $100$ to $9000$ comparisons. We also observe that the change in the win-rate is within $1\%$ when the dataset size increases from $4000$ to $9000$. This highlights that the performance gains are non-linear with the dataset size. In the future, it would be pertinent to explore techniques for selecting a subset of joint preference comparisons that result in maximum performance gains.
Figure 14: Results for scaling the feedback data size on TL;DR summarization dataset. We find that the win-rate improves with the increase in the dataset size using the Dove preference optimization objective.

### v2

@@
 A common technique for aligning large language models (LLMs) relies on acquiring human preferences by comparing multiple generations conditioned on a fixed context.
-This only leverages the pairwise comparisons when the generations are placed in an identical context.
-However, such conditional rankings often fail to capture the complex and multidimensional aspects of human preferences.
-In this work, we revisit the traditional paradigm of preference acquisition and propose a new axis that is based on eliciting preferences jointly over the instruction-response pairs.
-While prior preference optimizations are designed for conditional ranking protocols (e.g., DPO), our proposed preference acquisition protocol introduces Dove, a new preference optimization objective that upweights the joint probability of the chosen instruction-response pair over the rejected instruction-response pair.
-Interestingly, we find that the LLM trained with joint instruction-response preference data using Dove outperforms the LLM trained with DPO by $5.2\%$ and $3.3\%$ win-rate for the summarization and open-ended dialogue datasets, respectively.
+This method, however, relies solely on pairwise comparisons, where the generations are evaluated within an identical context.
+While effective to such conditional preferences often fail to encompass the nuanced and multidimensional nature of human preferences.
+In this work, we revisit the traditional paradigm of preference acquisition and propose a new axis based on eliciting preferences jointly over the instruction-response pairs.
+Unlike prior preference optimizations, which are designed for conditional ranking protocols (e.g., DPO), we propose Joint Preference Optimization (JPO), a new preference optimization objective that upweights the joint probability of the chosen instruction-response pair over the rejected instruction-response pair.
+Interestingly, LLMs trained with joint instruction-response preference data using JPO outperform LLM trained with DPO by $5.2\%$ and $3.3\%$ win-rate for summarization and open-ended dialogue datasets, respectively.
 Our findings reveal that joint preferences over instruction and response pairs can significantly enhance the alignment of LLMs by tapping into a broader spectrum of human preference elicitation.
 The data and code is available at https://github.com/Hritikbansal/dove.
+Comparing Bad Apples to Good Oranges:
+Aligning Large Language Models via Joint Preference Optimization
+Hritik Bansal♡∗      Ashima Suvarna♡∗      Gantavya Bhatt♢∗ Nanyun Peng     Kai-Wei Chang      Aditya Grover University of California, Los Angeles University of Washington, Seattle {hbansal,asuvarna31}@cs.ucla.edu,{gbhatt2}@uw.edu
 1 Introduction
-Recently, alignment [42, 31] has emerged as a crucial step in enhancing the performance of large language models (LLMs) [1, 30, 44, 3, 11, 48, 19] in diverse real-world applications [22, 65, 57, 12, 21].
-In particular, the aligned LLMs can produce high-quality outputs that satisfy human values and intents, including helpfulness, coherence, and harmlessness [4, 31].
-LLM alignment hinges on the quality of preferences acquired from human or AI annotators [31].
-Among the various preference acquisition protocols [23, 58, 39, 8], the ranking-based approach is the most widely used paradigm for aligning LLMs[42, 31, 6, 49, 45].
-Specifically, in this approach the annotator has to compare a pair of responses conditioned on a fixed context.
-For instance, human can select a ‘preferred’ response by comparing a pair of responses for the instruction ‘Create a list of four fruits other than Apple’ (Figure 1 (left)).
-Besides ranking preferences conditioned on a fixed context, humans can also express preferences in non-identical contexts.
-For example, while browsing reviews for products on an e-commerce website, humans are likely to prefer an accurate and detail-oriented review for a camera over an incoherent, vague movie review even though the products (camera and movie) are qualitatively different.
-Although the traditional conditional rankings provide rich preference for alignment, they fail to holistically capture the various dimensions of reasoning of human preferences.
-In this work, we revisit the traditional paradigm of conditional preference acquisition and propose a new approach for jointly eliciting preferences over instruction-response pairs.
-This method aims to uncover diverse reasoning paths in the process of acquiring feedback.
-In this work, we develop a framework to acquire preferences jointly over instruction-response pairs.
-Starting from an instruction-response data consisting of response $R_{i}$ for instruction $I_{i}$ (say $i\in\{1,2\}$), we acquire ranking-based preferences over the instruction-response pairs $(I_{1},R_{1})$ and $(I_{2},R_{2})$.
-As shown in Figure 1 (right), we aim to understand whether the response in the pair $X$ is perceived better than the response in the pair $Y$.
-For instance, humans would prefer a helpful response to the instruction ‘Create a list of four fruits’ over a response that completely ignores the instruction ‘Create a list of beach activities’.
-This suggests that we can reveal preference axes like adherence to instructions, grammatical fluency, and clarity, even when following a joint preference optimization protocol.
-In addition, our protocol can elicit human preference behaviours that are obfuscated in the prior protocols, and redefines conditional preference elicitation as a special case where the instructions are identical.
-Figure 1: Overview of the proposed Joint Preference Optimization paradigm.
+Recently, alignment (Stiennon et al., 2020; Ouyang et al., 2022) has emerged as a crucial step in enhancing the performance of large language models (LLMs) (Anthropic, 2024; OpenAI, 2023; Team et al., 2023; Anthrophic, 2023; Brown et al., 2020; Touvron et al., 2023; Jiang et al., 2023) in diverse real-world applications (Li et al., 2023; Zheng et al., 2023a; Wu et al., 2023a; Clusmann et al., 2023; Lambert et al., 2024).
+In particular, aligned LLMs generate responses that maximize human utility along various dimensions such as helpfulness, coherence, and harmlessness (Askell et al., 2021; Ouyang et al., 2022).
+Here, the notion of human utility is subjective Kirk et al.
+(2024); Gabriel (2020), and mainly hinges on how preferences are acquired from annotators Otto et al.
+(2022).
+Among the various preference acquisition protocols (Lightman et al., 2023; Wu et al., 2023b; Scheurer et al., 2023; Bansal et al., 2023), the ranking-based approach is the most widely used paradigm to align LLMs (Stiennon et al., 2020; Ouyang et al., 2022; Bai et al., 2022a; Tunstall et al., 2023; Teknium, 2023).
+Specifically, in ranking approach the annotator has to compare a pair of responses conditioned on a fixed context.
+For instance, humans can select a ‘preferred’ response by comparing a pair of responses for the instruction ‘Create a list of four fruits other than Apple’ (Figure 1 (left)).
+Figure 1: Overview of the Joint Preference Optimization.
 (Left) We show that the conditional preference acquisition method would require the annotators to compare two responses for an identical instruction.
@@
 Our framework thus elicits preferences that are obfuscated in the prior approach.
-Next, we propose Dove, a framework for aligning LLMs with our proposed joint preference elicitation scheme.
+Although traditional conditional rankings provide rich preferences for alignment, they fail to holistically capture the multi-faceted nature of human decision-making and preferences (Zhi-Xuan et al., 2024; Gigerenzer, 2008).
+Besides ranking preferences conditioned on a fixed context, humans can also express preferences in non-identical contexts.
+For example, while browsing reviews for products on an e-commerce website, humans are likely to prefer an accurate and detail-oriented review for a camera over an incoherent, vague movie review even though the products (camera and movie) are qualitatively different.
+In this work, we revisit the traditional paradigm of conditional preference acquisition by developing a framework to acquire preferences jointly over instruction-response pairs.
+Starting from an instruction-response data consisting of response $R_{i}$ for instruction $I_{i}$ (say $i\in\{1,2\}$), we acquire ranking-based preferences over the instruction-response pairs $(I_{1},R_{1})$ and $(I_{2},R_{2})$.
+As shown in Figure 1 (right), we aim to understand whether the response in the pair $X$ is perceived better than the response in the pair $Y$.
+We hypothesize that by capturing preferences in non-identical contexts our protocol can elicit human behaviors that are obfuscated in prior protocols.
+First, wee show that humans can provide decisive preferences in joint preferences protocol (5.4).
+Then, we analyze how joint preferences differ from conditional preferences on the same dataset(§5.4).
+Preference Acquisition Algorithm Alignment Objective Different Instructions
+Score Ethayarajh et al. (2024) Conditional No
+Comparison (DPO Variants) Rafailov et al. (2024) Conditional No
+Park et al. (2024) Conditional No
+Liu et al. (2024) Conditional No
+Meng et al. (2024) Conditional No
+Hong et al. (2024) Conditional No
+Pairwise JPO (ours) Joint Yes
+Table 1: JPO differs from prior works along three key aspects: preference acquisition (scoring or comparison), objective (conditional or joint distribution), and their ability to handle non-identical instruction-responses.
+Prior works like DPO and its variants Rafailov et al.
+(2023); Yin et al.
+(2024); Liu et al.
+(2024); Meng et al.
+(2024); Hong et al.
+(2024); Azar et al.
+(2023) rely on conditional rankings, and thus do not have access to the joint distribution of human preferences in the ranking protocol (Table 1).
+While a rating protocol Ethayarajh et al.
+(2024) allows for a comparison between responses from non-identical instructions, it can be inconsistent with rankings Bansal et al.
+(2023) and ignores the possibility of preferences over a pair of chosen or rejected responses.
+[^1] For instance, a pair of responses that achieves a score of 0, under the rating protocol, will result in an indecisive preference.
+Thus, we propose Joint Preference Optimization (JPO), a framework for aligning LLMs with our proposed joint preference elicitation scheme.
 Specifically, it upweights the joint probability of the chosen instruction-response pair over the rejected instruction-response pair.
-This differs from the other frameworks that assume conditional rankings in their feedback data, such as DPO [35, 5] and preference optimizations that train a separate reward model such as PPO and rejection sampling [40, 29].
-Further, we further point that Dove subsumes the prior preference optimizations as conditional rankings are a special case of joint preferences (e.g., when $I_{1}=I_{2}$).
-In our experiments, we focus on extending and comparing against DPO because of their simplicity, stability, and high-performance.
-However, our framework can be easily applied to PPO-like approaches by training a reward model on the joint preferences.
-Finally, we conduct experiments to explore the new reasoning paths enabled by joint preference elicitation, followed by aligning LLMs with the Dove objective.
-To do so, we explore the interplay between the feedback data collected under conditional rankings and joint preferences protocol.
-In addition, we ask human annotators to explain their preference decisions, uncovering new reasoning paths that highlight the complexities of the preference acquisition process (§4).
-After feedback acquisition, we aim to investigate the impact of diverse preferences collected from conditional and joint preferences on LLM alignment.
-In our experiments, we align a Mistral-7B LLM with the preferences acquired from the conditional rankings and joint preferences, using our Dove algorithm.
-We find that the Dove outperforms the supervised finetuned LLM by $30\%$ and $18\%$ win-rate against the gold responses on the unseen instructions from the summarization and open-ended dialogues datasets, respectively.
-Surprisingly, we find that Dove can effectively tap into the diverse preferences in the conditional and joint feedback data and outperforms DPO by $5.2\%$ and $3.3\%$ win-rate points on the summarization and open-ended dialogues, respectively.
+Furthermore, JPO subsumes prior preference optimizations as conditional rankings are a special case of joint preferences (e.g., when the instructions are identical).
+We conduct experiments to explore new reasoning paths enabled by joint preference elicitation and alignment of LLMs with the JPO objective.
+By analyzing feedback from conditional rankings and joint preferences protocols, along with explanations from human annotators, we uncover the complexities of the preference acquisition process (§5).
+Using our JPO algorithm, we align a Mistral-7B LLM with the collected preferences, achieving a $30\%$ and $18\%$ higher win rate against gold responses on unseen instructions from summarization and dialogue datasets, respectively.
+JPO leverages diverse preferences effectively, outperforming DPO by $5.2\%$ and $3.3\%$ win-rate points on the summarization and open-ended dialogues, respectively.
+In addition, JPO outperforms KTO by $3.5\%$ on the open-ended dialogues dataset.
+It also surpasses both DPO and KTO in the AlpacaEval2 benchmark (§6).
 This indicates that by utilizing the diverse preference signals present in the existing data, we can align an LLM robustly without acquiring additional instruction-response data.
+2 Related Work
+Alignment using Reinforcement Learning.
+Aligning LLMs with human preferences using reinforcement learning is widely adopted to ensure LLMs follow user intents without being harmful Ouyang et al.
+(2022).
+This alignment is usually done by first optimizing for a reward model on preference data Bradley and Terry (1952); Likert (1932); Bansal et al.
+(2023), followed by aligning the LLMs distribution that maximizes the learned reward model using Reinforcement Learning (RLHF) Schulman et al.
+(2017); Ouyang et al.
+(2022), with optional Divergence penalty Wang et al.
+(2023a) to avoid deviating from the reference policy.
+Additionally, Dubois et al.
+(2023); Lu et al.
+(2024); Zheng et al.
+(2023b) observe that preferences from LLMs can also be used for alignments motivating Reinforcement Learning through AI feedback (RLAIF).
+Contrary to prior work that collect preferences as conditional rankings, we emphasize that preference acquisition is a complex phenomenon and elicit joint preferences over instruction-response data.
+Reward Free Policy Alignment.
+Rafailov et al.
+(2024) introduced Direct Preference Optimization (DPO) that optimizes directly within the model parameter space, hence eliminating the reward modeling step.
+Liu et al.
+(2024) extends this framework where instead of two responses, alignment is done over the list of responses while Liu et al.
+(2023) improves DPO using statistical rejection sampling.
+Amini et al.
+(2024) provides an offset in the DPO objective to increase the margins and Pal et al.
+(2024) suggests adding an explicit penalty term to avoid a reduction in the likelihood of preferred pairs over the DPO training.
+Recent variants of DPO such as SimPO (Meng et al., 2024) alleviates the need of reference policy in the objective.
+Contrary to our work where we compare the joint distributions, Yin et al.
+(2024) proposes RPO that compares the conditional likelihood of a winning response with the losing response of another prompt.
+Beyond DPO, Ethayarajh et al.
+(2024) proposed a human-aware loss function-based framework using prospect theory named KTO, and Azar et al.
+(2023) proposes IPO that uses human preferences expressed as pairwise preferences.
+Lastly, Zhao et al.
+(2022) uses sequence likelihood calibration to align the model from human preference.
+Despite of a vast body of work arising from DPO, none of the existing methods can operate and contrast over the joint distribution of instruction-response pairs like the proposed JPO algorithm.
 3 Background
-In this work, our aim is to align language models to generate outputs that are preferred by humans across various dimensions such as helpfulness and coherence.
-The process of aligning a base model, which is pretrained on a large corpus of text [13, 38, 41, 33], involves multiple steps: (a) instruction-response data collection, (b) supervised fine-tuning, (c) preference data acquisition, and (d) deployment of an alignment algorithm.
-The instruction-response data can be either hand-crafted by humans [14, 54] or generated by machines [43, 49].
-Subsequently, the base model undergoes supervised fine-tuning (SFT) on the instruction-response pairs [66, 56, 54, 34, 59, 17, 60, 55, 62, 47].
-Following SFT, feedback data is acquired under a specific acquisition protocol (e.g., rankings) from the annotators (§2.1).
-Finally, an alignment algorithm trains the SFT model on the feedback data (§2.2).
-2.1 Ranking Feedback Acquisition Protocol
-Assume a supervised finetuned language model $p_{\text{sft}}$ that is capable of responding to user instructions (e.g., imperative tasks or questions).
-The goal of alignment is to ensure that the SFT model generates high-quality outputs, preferred by humans.
-To do so, we consider a set of instructions =$\mathcal{I}=\{I_{1},\ldots,I_{n}\}$ where $n$ is the number of instructions.
-Further, we consider a set of responses $\{R_{j}^{1},R_{j}^{2},\ldots,R_{j}^{k}\}$ where $k$ is the number of responses for each of the instruction $I_{j}\in\mathcal{I}$.
-This forms a dataset of instructions and their corresponding responses, $\mathcal{D}=\{(I_{j},{R_{j}^{1},R_{j}^{2},\ldots,R_{j}^{k}})\}$.
-[^1] We will drop the iterator over j when defining the dataset for the ease of notation.
-Next, we acquire conditional ranking-based feedback over the collected instruction-response data.
-Under this feedback acquisition protocol, the annotator selects a chosen and rejected response from $\{R_{j}^{x},R_{j}^{y}\}$ conditioned on the instruction $I_{j}$ where $x,y\in\{1,2,\ldots,k\}$.
-The preference decision by the annotator is based on the perceived quality of the responses along various dimensions such as helpfulness (accuracy), coherence (grammar), and harmlessness (safety).
-Formally, the annotator assigns a instruction-conditioned ranking feedback $c(I_{j},R_{j}^{x},R_{j}^{y})\in\{R_{j}^{x},R_{j}^{y},\text{Equal}\}$ where ‘Equal’ indicates that both responses are perceived equally good or bad.
-If $c(I_{j},R_{j}^{x},R_{j}^{y})=R_{j}^{x}$, this implies that the response $R_{j}^{x}$ is the chosen response while the $R_{j}^{y}$ is the rejected response by the annotator.
-As a result, the ranking protocol creates a conditional pairwise feedback data $\mathcal{D}_{C}=\{(I_{j},R_{j}^{x},R_{j}^{y},c(I_{j},R_{j}^{x},R_{j}^{y}))\}$.
-Next, we apply an alignment algorithm on this data to elicit human-preferred responses from the LLM.
-2.2 Alignment Algorithm
-Rafailov et al.
-[36] introduced direct preference optimization (DPO) that can align a language model without utilizing on an external reward model.
-Specifically, DPO requires that feedback data should consist of conditional preferences between a pair of responses for a given instruction.
-Additionally, the algorithm assumes a preference dataset $\mathcal{D}_{C}$ and the reference model $p_{\text{ref}}$ which is usually the supervised finetuned language model $p_{\text{sft}}$.
-Specifically, it aims to train an aligned model $p_{\theta}$ using an optimization objective that upweights the conditional probability of the chosen response $p_{\theta}(R_{j}^{w}|I_{j})$ over the rejected response $p_{\theta}(R_{j}^{\ell}|I_{j})$ where $R_{j}^{w}$ and $R_{j}^{\ell}$ are the chosen and rejected response, respectively.
-Formally, the optimization objective for DPO minimizes:
-$\mathcal{L}_{\textsc{DPO}}(\theta;\mathcal{D}_{C},\beta,p_{\text{ref}})={\mathbb{E}}_{(I_{j},R_{j}^{w},R_{j}^{\ell})\sim\mathcal{D}_{C}}\left[\operatorname{log}\left(\sigma\left(\beta\operatorname{log}\frac{p_{\theta}(R_{j}^{w}|I_{j})}{p_{\text{ref}}(R_{j}^{w}|I_{j})}-\beta\operatorname{log}\frac{p_{\theta}(R_{j}^{\ell}|I_{j})}{p_{\text{ref}}(R_{j}^{\ell}|I_{j})}\right)\right)\right]$ (1)
-where $\sigma$ denotes the sigmoid function and $\beta$ is a hyperparameter. Post-alignment, the model generates high-quality outputs for unseen instructions that are preferred by the annotators.
-3 Joint Preference Optimization using Dove
-3.1 Joint Preference Acquisition Protocol
-In §2.1, we describe a common technique for feedback data acquisition that requires the annotators to assign a preferred and non-preferred label to a pair of responses for an instruction.
-However, this paradigm does not capture the complex and multidimensional aspects of human preferences [20, 46].
-Specifically, the reasoning paths for making preference decisions depend upon the context in which the comparison is made.
+In this work, we focus on aligning language models to generate outputs preferred by humans in dimensions like helpfulness and coherence.
+Aligning a pretrained base model involves four steps: (a) collecting instruction-response data, (b) supervised fine-tuning (SFT), (c) acquiring preference data, and (d) deploying an alignment algorithm.
+The instruction-response data can be either hand-crafted by humans (Conover et al., 2023; Wang et al., 2022) or generated by AI (Taori et al., 2023; Tunstall et al., 2023).
+Subsequently, the base model undergoes supervised fine-tuning (SFT) on the instruction-response pairs (Zheng et al., 2023b; Wang et al., 2023c, 2022; Peng et al., 2023; Xu et al., 2023; Yin et al., 2023; Wang et al., 2023b; Yu et al., 2023; Toshniwal et al., 2024).
+Following SFT, feedback data is gathered (e.g., rankings) to train the SFT model via an alignment algorithm.
+This often involves training a reward model on preference data (Bradley and Terry, 1952; Bansal et al., 2023) and aligning the model using Reinforcement Learning (RLHF) (Schulman et al., 2017; Ouyang et al., 2022).
+To address challenges in human feedback collection (Dubois et al., 2023; Zheng et al., 2023b), LLMs can provide feedback, enabling Reinforcement Learning through AI Feedback (RLAIF).
+Alternatively, Rafailov et al.
+(2024) introduced Direct Preference Optimization (DPO) that mitigates the instability due to PPO for reward maximization by optimizing directly within the model parameter space, hence by-passing the reward modeling step.
+4 Joint Preference Optimization (JPO)
+A common technique for feedback data acquisition requires the annotators to assign a preferred and non-preferred label to a pair of responses for an instruction (Stiennon et al., 2020; Rafailov et al., 2023; Ouyang et al., 2022; Ethayarajh et al., 2024).
+However, this paradigm does not capture the complex and multidimensional aspects of human preferences (Kendall and Smith, 1940; Thurstone, 2017; Zhi-Xuan et al., 2024).
+Specifically, the heuristics for making preference decisions depend upon the context in which the comparison is made Otto et al.
+(2022).
 While the traditional ranking protocol compares the two responses under a fixed context, humans can perform pairwise comparisons jointly over instruction-response pairs.
 For example, consider two summaries, A and B, for articles X and Y, respectively; then, a human can reason and choose the response that better summarizes its corresponding article.
-Hence, it is critical to align language models with diverse feedback signals to elicit high-quality responses that humans prefer under various contexts.
+Hence, it is critical to align language models with diverse feedback signals to accurately model human behavior and decision making.
 In our setup, the annotator has to decide a chosen and rejected instruction-response pair $(I_{a},R_{a},I_{b},R_{b})$ where $R_{a}$ and $R_{b}$ are responses to the instructions $I_{a}$ and $I_{b}$, respectively, and $(I_{a},R_{a}),(I_{b},R_{b})\in\mathcal{D}$.
@@
 Finally, we assess the interplay between the joint feedback dataset $\mathcal{D}_{H}$ with the conditional feedback dataset $\mathcal{D}_{C}$ along with qualitative examples in §5.
-3.2 Dove
-Here, we propose Dove, a preference optimization objective that learns to align the language models with the preferences acquired jointly over the instruction-response pairs.
+We propose JPO, a preference optimization objective that learns to align the language models with preferences acquired jointly over the instruction-response pairs.
 We assume a joint preference dataset $\mathcal{D}_{X}=\{(I_{i}^{w},R_{i}^{w},I_{j}^{\ell},R_{j}^{\ell})\}$, that can be constructed from $\mathcal{D}_{H}$, where $(I_{i}^{w},R_{i}^{w})$ and $(I_{j}^{\ell},R_{j}^{\ell})$ are the chosen and rejected instruction-response pairs, respectively.
 Similar to DPO, we start with a reference model $p_{\text{ref}}$ which is usually the supervised finetuned language model $p_{\text{sft}}$.
-Specifically, the Dove objective aims to learn an aligned model $p_{\theta}$ by upweighting the joint probability of preferred responses $p(R_{i}^{w},I_{i}^{w})$ over non-preferred responses $p(R_{j}^{\ell},I_{j}^{\ell})$.
-Formally, the optimization objective for Dove minimizes:
-$\mathcal{L}(\theta;\mathcal{D}_{X},\beta,p_{\text{ref}})$ $=\mathbb{E}_{(I_{j}^{w},R_{j}^{w},I_{j}^{\ell},R_{j}^{\ell})\sim\mathcal{D}_{X}}\left[\log\left(\sigma\left(\beta\log\frac{p_{\theta}(R_{i}^{w},I_{i}^{w})}{p_{\text{ref}}(R_{i}^{w},I_{i}^{w})}-\beta\log\frac{p_{\theta}(R_{j}^{\ell},I_{j}^{\ell})}{p_{\text{ref}}(R_{j}^{\ell},I_{j}^{\ell})}\right)\right)\right]$ (2)
+Specifically, the JPO objective aims to learn an aligned model $p_{\theta}$ by upweighting the joint probability of preferred responses $p(R_{i}^{w},I_{i}^{w})$ over non-preferred responses $p(R_{j}^{\ell},I_{j}^{\ell})$.
+Formally, the optimization objective for JPO, $\mathcal{L}(\theta;\mathcal{D}_{X},\beta,p_{\text{ref}})$ minimizes the expectation over $(I_{j}^{w},R_{j}^{w},I_{j}^{\ell},R_{j}^{\ell})\sim\mathcal{D}_{X}$ :
+$\mathbb{E}\left[\log\left(\sigma\left(\beta\log\frac{p_{\theta}(R_{i}^{w},I_{i}^{w})}{p_{\text{ref}}(R_{i}^{w},I_{i}^{w})}-\beta\log\frac{p_{\theta}(R_{j}^{\ell},I_{j}^{\ell})}{p_{\text{ref}}(R_{j}^{\ell},I_{j}^{\ell})}\right)\right)\right]$ (1)
 where $\sigma$ denotes the sigmoid function and $\beta$ is a hyperparameter.
 Further, we show that Eq.
-3 reduces to the DPO formulation (Eq.
-1) when the instructions $I_{i}=I_{j}$ in Appendix §D.
-We can also see that the Dove objective aims to learn an aligned model $p_{\theta}$ by upweighting the conditional probability of preferred responses $p(R_{i}^{w}|I_{i}^{w})$ over non-preferred responses $p(R_{j}^{\ell}|I_{j}^{\ell})$, along with a correction factor based on the prior probability of the instructions under the language model $p_{\theta}(I_{i}^{w})$ and $p_{\theta}(I_{j}^{\ell})$.
-In §5, we utilize Dove to align language models to generate human-preferred summaries and answer open-ended instructions.
+1 reduces to the DPO formulation (Appendix Eq.
+2) when the instructions $I_{i}=I_{j}$ in Appendix §D.
+We can also see that the JPO objective aims to learn an aligned model $p_{\theta}$ by upweighting the conditional probability of preferred responses $p(R_{i}^{w}|I_{i}^{w})$ over non-preferred responses $p(R_{j}^{\ell}|I_{j}^{\ell})$, along with a correction factor based on the prior probability of the instructions under the language model $p_{\theta}(I_{i}^{w})$ and $p_{\theta}(I_{j}^{\ell})$.
+In §6, we utilize JPO to align language models to generate human-preferred summaries and answer open-ended instructions.
+4.1 Comparison of Joint Preferences with Prior Preference Protocols
+JPO improves over prior work by acquiring ranking-based preferences over non-identical instructions that has remained unexplored in prior work (please refer to table 1).
+Diverse human reasoning cannot be captured in the traditional conditional framework it fails to capture human preferences over varied contexts.
+Context influences decision-making and subjective valuation when capturing human preferences (Otto et al., 2022).
+Prior work Yin et al.
+(2024); Liu et al.
+(2024); Meng et al.
+(2024); Hong et al.
+(2024) collect conditional preferences in a pairwise manner and are variants of DPO Rafailov et al.
+(2023).
+Thus, in our experiments we compare JPO to DPO directly.
+Furthermore, we implement KTO Ethayarajh et al.
+(2024) as a baseline since KTO removes the requirements of preference data that should be paired in preference optimization and implicitly compares responses from different instructions.
+We find that JPO outperforms both DPO and KTO.
 5 Interplay between Feedback Protocols
 5.1 Instruction-Response Acquisition
-The instruction-response data is a collection of real-world queries that are presented to the text AI assistants.
 In this work, we consider two kinds of instruction-response data.
-First, we consider a filtered version of the TL;DR summarization dataset [51] from [42] consisting of Reddit posts, their summarizes, and human preferences over a pair of summaries for a given post.
+First, we consider a filtered version of the TL;DR summarization dataset (Völske et al., 2017) from Stiennon et al.
+(2020) consisting of Reddit posts, their summarizes, and human preferences over a pair of summaries for a given post.
 Throughout the dataset, the task is of summarization that is close-ended and well-defined for language models.
-Second, we consider the single-turn dialogues from the helpful-base subset of the Anthropic-HH dataset [7].
-Specifically, this dataset consists of open-ended instructions with a collection of responses ranging from ‘Which coffee bean is better for a morning roast?’ to ‘How do I attract more hummingbirds in my yard?’.
-Both these datasets come with the train and test split where each instance consists of an instruction and a pair of responses $\mathcal{D}=\{(I_{i},R_{i}^{1},R_{i}^{2})\}_{i=1}^{n}$ where $n$ is the dataset size.
-In this work, we collect AI and human feedback on the instruction-response data from their train split and filter instances where instructions are repeated.
+Second, we consider the single-turn dialogues from the helpful-base subset of the Anthropic-HH dataset (Bai et al., 2022b).
+Specifically, this dataset consists of open-ended instructions with a collection of responses.
+Both these datasets have a train and test split where each instance consists of an instruction and a pair of responses $\mathcal{D}=\{(I_{i},R_{i}^{1},R_{i}^{2})\}_{i=1}^{n}$ where $n$ is the dataset size.
+In this work, we collect AI and human feedback on the instruction-response data from their train split and filter instances with duplicate instructions.
 We can directly compare the two responses for the fixed instruction and construct a ranking feedback dataset $\mathcal{D}_{C}=\{(I_{i},R_{i}^{1},R_{i}^{2},c(I_{i},R_{i}^{1},R_{i}^{2}))\}$.
-To acquire preferences jointly over the instruction-response pairs, we select one of the responses, at random, from every instance of $\mathcal{D}$ to construct $\mathcal{D}_{S}=\{(I_{i},R_{i})\}$ where $R_{i}\in\{R_{i}^{1},R_{i}^{2}\}$.
+To acquire preferences jointly over the instruction-response pairs, we randomly select one of the responses from every instance of $\mathcal{D}$ to construct $\mathcal{D}_{S}=\{(I_{i},R_{i})\}$ where $R_{i}\in\{R_{i}^{1},R_{i}^{2}\}$.
 Subsequently, we create the joint instruction-response pairs by matching every instance $(I_{i},R_{i})\in\mathcal{D}_{S}$ with another instance $(I_{j},R_{j})\in\mathcal{D}_{S}$ to get $\mathcal{D}_{H}=\{(I_{i},R_{i},I_{j},R_{j},h(I_{i},R_{i},I_{j},R_{j}))\}$ of the same size as $\mathcal{D}_{S}$ and $\mathcal{D}_{C}$.
 In §6, we will utilize $\mathcal{D}_{S}$ to SFT the base model, and $\mathcal{D}_{C}$ and $\mathcal{D}_{H}$ as preference datasets for LLM alignment.
-We provide the dataset statistics in Appendix §E.
+We provide the dataset statistics in Appendix §C.
 5.2 Feedback from AI and Humans
+Dataset Ranking H-H H-AI
+TL;DR Conditional 69% 63%
+Anth.-Helpful 70.1% 72%
+TL;DR Joint (Non-Identical) 62% 60%
+Anth.-Helpful 68.8% 71%
+Average 67.5% 66.5%
+Table 2: Agreement analysis between within human annotators and gold human feedback and AI (ChatGPT) feedback.
+We perform the agreement calculations for the two ranking protocols: (a) conditional rankings, and (b) joint preferences where instructions are non-identical.
+In addition, we assess the agreement rates over the two datasets: (a) TL;DR and (b) Anthropic-helpful dataset.
 Feedback from AI.
-Prior work [15, 7] has shown that AI feedback can be leveraged to align language models to generate helpful and harmless responses to unseen instructions.
-In addition, acquiring AI feedback at large-scale is more accessible and cheaper in comparison to human feedback.
-To this end, we collect feedback over a pair of responses for a fixed instruction, and joint instruction-response pairs without identical instructions from GPT-3.5-Turbo-0125 (ChatGPT).
-The choice of ChatGPT was motivated by its affordability (e.g., output tokens from ChatGPT are 50×50$\times$50 × cheaper than GPT-4).
-To collect ranking feedback over a pair of responses for a fixed instruction, we prompt ChatGPT to assign a chosen response.
-To mitigate any bias from the ordering of the two responses, we run two queries for all comparisons to cover all possible orderings.
-When the ChatGPT preferences flip by flipping the order of the two responses, then we consider it a tie, similar to [8, 9].
-Specifically, the AI is instructed to provide its preference based on the accuracy, coherence, and harmlessness of the responses.
+We collect feedback over a pair of responses for a fixed instruction, and joint instruction-response pairs without identical instructions from GPT-3.5-Turbo-0125 (ChatGPT).
+We choose ChatGPT due to its affordability (e.g., output tokens from ChatGPT are 50×50$\times$50 × cheaper than GPT-4).
+To mitigate any ordering bias, we run two queries for all comparisons.
+When the ChatGPT preferences flip by flipping the order of the two responses, then we consider it a tie, similar to (Bansal et al., 2023; Bitton et al., 2023).
+Specifically, we instruct the AI to to choose the response that is more accurate, coherent, and harmless.
+To collect conditional preferences over a pair of responses for a fixed instruction, we prompt ChatGPT to choose a response.
 To collect AI preferences jointly over the instruction-response pairs, we prompt ChatGPT to decide the response that better answers its corresponding instruction.
-Similar to the previous scenario, we run two queries for all comparisons to mitigate any ordering bias and provide guidelines to choose the response that is more accurate, coherent, and harmless.
 We collected approximately $50K$ comparisons across both feedback acquisition protocols for the summarization and Anthropic-Helpful dataset, at a cost of \$$\$100$.
-We provide the AI prompts in Appendix §F.
+We provide the AI prompts in Appendix §H.
 Feedback from Humans.
-In this work, we also collect human preferences for $2000$ comparisons over summarization and Anthropic-Helpful dataset.
-Such a data is useful for providing insights into the human behavior under different preference acquisition protocols (§4.3).
-In addition, this data aids in agreement between the ChatGPT and human decisions.
-Specifically, we ask two annotators to assign a chosen response or choose ‘equal’ after comparing the quality of the responses along the same dimensions as ChatGPT guidelines.
-The human annotations were collected from Amazon Mechanical Turk (AMT) from the participants that passed a preliminary qualification exam.
+In this work, we also collect human preferences for $2000$ comparisons over TL;DR and Anthropic-Helpful dataset.
+Specifically, we ask two annotators to assign a chosen response or chosen instruction-response pair based along the same dimensions as ChatGPT guidelines.
+Annotators can also choose ‘equal’ if they fail to make a identify a decisive preference.
+The human annotations were collected from Amazon Mechanical Turk (AMT).
+We recruited the participants that passed a preliminary qualification exam.
 In total, we spent \$$\$720$ on human feedback acquisition.
-We provide the screenshot of the annotation UI in Appendix §G.
-We find that the AI achieves an agreement score of $64\%$ with the human annotators in our agreement analysis, compared to the $69\%$ inter-annotator agreement (Section C).
-This highlights the high quality of the AI feedback across datasets and preference protocols (i.e., conditional and joint preferences over non-identical instructions).
+We provide the screenshot of the annotation UI in Appendix §I.
+5.3 Agreement Analysis
+Figure 2: Results for the preferences acquired jointly over the instruction-response pairs where both the responses were either chosen or rejected under the conditional rankings protocol.
+Here, decisive implies that the annotators could assign a preference to one instruction-response pair over the other.
+Here, AH means Anthropic-Helpful.
+Figure 3: Results for the preferences acquired jointly over the instruction-response pairs where one of the instruction-response pair was chosen (C) and the other pair was rejected (R) under the conditional rankings.
+Here, $C<R$ implies that the instruction-response pair that was rejected under conditional rankings is actually preferred over an instruction-response pair that was rejected under the conditional rankings.
+Here, AH means Anthropic-Helpful.
+We present the annotator agreement scores in Table 2.
+We find that the average agreement is $67.5\%$ and $66.5\%$ between the human-human and human-AI annotators, respectively.
+Specifically, we find that in the conditional setup (identical instructions), the average human-human agreement is 69.5% for the TLDR and Anthropic-Helpfulness datasets.
+Similarly, in the joint setup (non-identical instruction-response pairs), the average inter-rater agreement is 68% on the same datasets.These agreement scores are comparable to those reported in prior studies (Li et al., 2023; Bansal et al., 2023), demonstrating the robustness of our evaluation.
+Interestingly, we find that agreement scores vary depending on the underlying distribution of instruction-response pairs and the choice of ranking protocol.
+Overall, our results highlight that humans and AI can provide rich feedback in both conditional and joint setup with acceptable agreement.
+Table 3: Response B and D are rejected under the conditional rankings.
+However, in the the joint setup humans consider that the response B answers Instruction 1 better than response D answers Instruction 2.
+Thus, joint setup human give a decisive feedback between two responses that were rejected under the conditional rankings.
+Instruction 1 My dog has been scratching a lot, but he doesn’t have any fleas. What else could be causing this?
+Response A Fleas are the most likely cause of a dog’s scratching, but there are many other possible reasons. It’s worth taking him to the vet to check things out.
+Response B Oh no! Maybe his skin is dry. I bet if you gave him a good bath, that would help!
+Conditional Preference 1 Human Preference: A>B. Human Explanation: Response A answered the query by giving possible causes of the problem along with a suggestion of just going to the vet.
+Instruction 2 Do you have any good chili recipes?
+Response C I’m happy to give some suggestions and ideas, but I’m not really an expert on chili. What kind of chili do you want to make?
+Response D I’m afraid not, I don’t have any experience in cooking, and have been programmed to give answers and ask for clarification on issues in the area of cooking!
+Conditional Preference 2 Human Preference: C>D.
+Human Explanation: Response D claims they can’t help and then later on in the same sentence claims they’ve been programmed to give answers on cooking issues, which doesn’t make much sense.
+C asks a reasonable clarifying question.
+Joint preference Human Preference: (Instruction 1, Response B) > (Instruction 2, Response D).
+Human Explanation: Response B provides answer to the Instruction 1 while Response D does not provide an accurate answer to Instruction 2.
 5.4 Interplay Analysis
 Setup.
-Here, we aim study the interaction between the conditional rankings and joint rankings over non-identical instructions.
+Here, we aim to study the interaction between the conditional rankings and joint rankings over non-identical instructions.
 Formally, each instruction-response pair $(I_{i},R_{i}^{x})$ from the conditional pairwise feedback dataset $\mathcal{D}_{C}$ where $x\in\{1,2\}$ can be assigned a preference $\mathcal{P}_{C}(I_{i},R_{i}^{x})$ among {‘chosen’, ‘reject’, ‘equal’}.
-For instance, $\mathcal{P}_{C}(I_{i},R_{i}^{1})=`\text{chosen}^{\prime}$ and $\mathcal{P}_{C}(I_{i},R_{i}^{2})=`\text{reject}$ if the response $R_{i}^{2}$ is rejected in the dataset $\mathcal{D}_{C}$ i.e., $c(I_{i},R_{i}^{1},R_{i}^{2})=R_{i}^{1}$.
+For instance, $\mathcal{P}_{C}(I_{i},R_{i}^{1})=`\text{chosen}^{\prime}$ and $\mathcal{P}_{C}(I_{i},R_{i}^{2})=`\text{reject}^{\prime}$ if the response $R_{i}^{2}$ is rejected in the dataset $\mathcal{D}_{C}$ i.e., $c(I_{i},R_{i}^{1},R_{i}^{2})=R_{i}^{1}$.
 Similarly, we can assign a preference $\mathcal{P}_{H}(I_{i},R_{i})$ among {‘chosen’, ‘reject’, ‘equal’} to an instruction-response pair $(I_{i},R_{i})$ from the joint preference dataset $\mathcal{D}_{H}$.
-For instance, $\mathcal{P}_{H}(I_{i},R_{i})=`\text{chosen}^{\prime}$ and $\mathcal{P}_{H}(I_{j},R_{j})$ where $i!=j$ if the instruction-response pair $(I_{i},R_{i})$ is chosen in the dataset $\mathcal{D}_{H}$ i.e., $h(I_{i},R_{i},I_{j},R_{j})=R_{i}$.
+For instance, $\mathcal{P}_{H}(I_{i},R_{i})=`\text{chosen}^{\prime}$ and $\mathcal{P}_{H}(I_{j},R_{j})=`\text{reject}^{\prime}$ where $i!=j$ if the instruction-response pair $(I_{i},R_{i})$ is chosen in the dataset $\mathcal{D}_{H}$ i.e., $h(I_{i},R_{i},I_{j},R_{j})=(I_{i},R_{i})$.
 To study the interplay between the preference protocols, we assess $\mathcal{P}_{C}(I_{i},R_{i})$, $\mathcal{P}_{C}(I_{j},R_{j})$, $\mathcal{P}_{H}(I_{i},R_{i})$ and $\mathcal{P}_{H}(I_{j},R_{j})$ for all $(I_{i},R_{i},I_{j},R_{j})\in\mathcal{D}_{H}$.
-Here, if $\mathcal{P}_{H}(I_{i},R_{i})=`\text{chosen}^{\prime}$ then $\mathcal{P}_{H}(I_{i},R_{i})=`\text{reject}^{\prime}$.
-For instance, if $\mathcal{P}_{C}(I_{i},R_{i})=`\text{chosen}^{\prime}$ and $\mathcal{P}_{C}(I_{j},R_{j})=`\text{chosen}^{\prime}$ then it implies that the annotators can reason about the joint preferences over a pair of instruction-response pairs that are originally preferred under the conditional ranking feedback protocol.
-We quantify study the interplay between the two ranking-based feedback from AI and Human annotators over summarization and open-ended Anthropic-Helpful datasets.
-Results.
-Data (Annotator) Decisive Indecisive
-TL;DR (AI) 63.7% 36.2%
-TL;DR (Human) 73.8% 25.7%
-Anthropic-Helpful (AI) 68.5% 31.5%
-Anthropic-Helpful (Human) 77.9% 22.0%
-Average 71.0% 29.0%
-(a) Results for the preferences acquired jointly over the instruction-response pairs where both the responses were either chosen or rejected under the conditional rankings protocol.
-Here, decisive implies that the annotators could assign a preference to one instruction-response pair over the other.
-Data (Annotator) C $>$ R C $<$ R Indecisive
-TL;DR (AI) 53.3% 14.3% 30.4%
-TL;DR (Human) 41.6% 22.2% 36.1%
-Anthropic-Helpful (AI) 54.5% 17.6% 27.8%
-Anthropic-Helpful (Human) 57.1% 21.4% 21.4%
-Average 52.0% 19.0% 29.0%
-(b) Results for the preferences acquired jointly over the instruction-response pairs where one of the instruction-response pair was chosen (C) and the other pair was rejected (R) under the conditional rankings.
-Here, $C<R$ implies that the instruction-response pair that was rejected under conditional rankings is actually preferred over an instruction-response pair that was rejected under the conditional rankings.
-Table 1: Results for the interplay between the conditional rankings and joint rankings acquired over non-identical instructions for summarization and open-ended dialogue dataset.
-In total, we compare $48K$ and $1K$ annotations from the AI and humans, respectively.
-We present the results for the interaction analysis in Table 1.
-In Table 1(a), we study the joint preferences over the instruction-response pairs $(I_{i},R_{i},I_{j},R_{j})$ where the individual instruction and response data is either chosen or rejected in the conditional feedback protocol (e.g., $\mathcal{P}_{C}(I_{z},R_{z})=`\text{chosen}^{\prime}$ for $z\in\{i,j\}$).
+Here, if $\mathcal{P}_{H}(I_{i},R_{i})=`\text{chosen}^{\prime}$ then $\mathcal{P}_{H}(I_{j},R_{j})=`\text{reject}^{\prime}$.
+Annotators show decisiveness in joint setup.
+In Figure 2, we study the joint preferences over the instruction-response pairs $(I_{i},R_{i},I_{j},R_{j})$ where the individual instruction and response data is either chosen or rejected in the conditional feedback protocol (e.g., $\mathcal{P}_{C}(I_{z},R_{z})=`\text{chosen}^{\prime}$ for $z\in\{i,j\}$).
 Interestingly, we find that the annotators can assign a decisive preference (e.g., $(I_{i},R_{i})>(I_{j},R_{j})$) in $71\%$ of the joint comparisons.
@@
 This highlights the existence of valid preference decisions that remained obfuscated in the traditional approach for ranking-based feedback acquisition.
-In Table 1(b), we study the joint preference over the instruction-response pairs $(I_{i},R_{i},I_{j},R_{j})$ where one of them is chosen and the other is rejected in the conditional feedback protocol (e.g., $\mathcal{P}_{C}(I_{i},R_{i})=`\text{chosen}^{\prime}$ and $\mathcal{P}_{C}(I_{i},R_{i})=`\text{reject}^{\prime}$).
+Annotator preferences depend on context and comparisons.
+In Figure 3, we study the joint preference over the instruction-response pairs $(I_{i},R_{i},I_{j},R_{j})$ where one of them is chosen and the other is rejected in the conditional feedback protocol (e.g., $\mathcal{P}_{C}(I_{i},R_{i})=`\text{chosen}^{\prime}$ and $\mathcal{P}_{C}(I_{j},R_{j})=`\text{reject}^{\prime}$).
 To our surprise, we find that the annotators do not prefer the instruction-response pair that was chosen under the conditional feedback protocol in $48\%$ of the comparisons.
@@
 This highlights that both human and AI annotators’ perceptions of preferred and non-preferred data depends on the context of the comparisons, indicating that feedback acquisition is a multifaceted phenomenon.
-Qualitative Examples.
-To probe the reasoning paths of the human annotators used for decision making, we ask them to provide brief explanations for their feedback decisions regarding a few conditional and joint preferences.
-We provide a list of qualitative examples consisting of instructions, responses, and respective preferences in Appendix §H.
-In Figure 8, we discovered that human annotators provided decisive feedback when comparing instruction-response pairs, basing their decisions on the accuracy of the responses.
-In Figure 11, we find that the human annotators preferred a instruction-summary pair, that was rejected under the conditional preference, because it provides a fuller picture of the original reddit post.
-In summary, we expose the multi-faceted reasoning paths of humans in joint instruction-response feedback acquisition that would have been concealed in the conditional feedback acquisition paradigm.
-5 LLM Alignment
-In the previous sections, we show that the humans and AI are capable of providing ranking-based feedback for a pair of responses for identical and non-identical instructions in the context.
-Here, we aim to study the impact of feedback data on eliciting high-quality responses from the large language models.
+Qualitative Case Study.
+To understand the heuristics used in preference annotations, we asked human annotators to provide brief explanations for their feedback decisions in both conditional and joint preference setups.
+In Table 3, we observe that humans provide reasonable explanations for rejecting responses B and D in the conditional setup.
+However, when these same rejected responses are presented in a joint setup, humans offer decisive feedback, basing their decisions on the accuracy of the responses—an aspect not emphasized in the explanations for the conditional preferences.
+We present additional qualitative examples in Appendix §F to showcase the multi-faceted nature of human feedback revealed through joint preferences.
+6 LLM Alignment with JPO
+(a) Performance on TL;DR
+(b) Performance on Anthropic-Helpful
+(c) Performance on AlpacaEval2.0
+Figure 4: Results for aligning LLMs with JPO.
+We utilize ChatGPT to compare the model responses with the gold responses.
+In 4a and 4b we report the results averaged over three runs of the preference optimization objectives and three sampling temperatures.
+In 4c, we report the results for temeperature set at 0.7 for AlpacaEval2.
+In previous sections, we explore how we collect ranking-based feedback for a pair of responses for identical and non-identical instructions.
+Here, we study how to leverage joint and conditional feedback data to align large language models effectively with JPO §4.
 6.1 Setup
-Here, we aim to align Mistral-7B  [19], a strong base LLM for its model capacity.
+We align Mistral-7B  (Jiang et al., 2023), a strong base LLM for its model capacity.
 We experiment with two datasets that exhibit diverse characteristics: (a) TL;DR dataset where the instruction is to summarize Reddit posts, and (b) open-ended dialogues from Anthropic-Helpful dataset (§5.1).
-In particular, we collect a conditional preference data $\mathcal{D}_{C}$ and joint preference data for non-identical instructions $\mathcal{D}_{H}$ of similar data sizes on both the datasets from ChatGPT.
+In particular, we collect a conditional preference data $\mathcal{D}_{C}$ and joint preference data for non-identical instructions $\mathcal{D}_{H}$ of similar data sizes from ChatGPT.
 Then, we convert the conditional preference data into an instruction-response data for supervised finetuning $\mathcal{D}_{\text{SFT}}$.
-First, we supervise finetune the entire base LLM model parameters with the SFT dataset to ensure that the preference data is in-policy for the alignment algorithms  [36].
-Subsequently, we apply DPO algorithm on the SFT model using the conditional preference data for $10$ epochs and $5$ epochs for the summarization and Anthropic-helpful data, respectively.
-Specifically, we use low-rank adaptation  [18] of SFT model during DPO alignment.
-The DPO optimization was trained on a single GPU Nvidia A6000 with a batch size of $32$.
-We note that our proposed Dove algorithm can utilize both the conditional preferences and joint preference with non-identical context.
-It is because the conditional preferences can be viewed as joint preferences with identical context.
-As a result, we train the base LLM with Dove algorithm after merging conditional and joint preferences data $\mathcal{D}_{M}=\mathcal{D}_{C}\cup\mathcal{D}_{H}$.
-We keep the hyperparameters (e.g., $\beta$), number of epochs, and the batch size identical to the DPO algorithm.
-In our experiments, we also train Dove algorithm on the joint preferences with non-identical instructions and highlight their usefulness for LLM alignment.
-We provide more details on training setup in Appendix §I.
+We supervise finetune the entire base LLM model parameters with the SFT dataset to ensure that the preference data is in-policy for the alignment algorithms  (Rafailov et al., 2023).
+JPO algorithm can utilize both the conditional preferences and joint preference with non-identical context.
+[^3] It is because the conditional preferences can be viewed as joint preferences with identical context.
+Thus, we train the base LLM with JPO algorithm after merging conditional and joint preferences data $\mathcal{D}_{M}=\mathcal{D}_{C}\cup\mathcal{D}_{H}$.
+We provide more details on training setup in Appendix §G.
+We also apply we apply DPO and KTO algorithm on the SFT model to compare against JPO.
 Post-alignment, we evaluate the aligned model responses against the gold responses in the dataset’s test split.
-Specifically, both datasets come with a human-preferred response for an instruction, which is treated as the gold response.
 We utilize ChatGPT to compare model and gold responses to decide on the preferred response or a tie.
-Finally, we report the win-rate of the model responses as the evaluation metric for $500$ unseen instructions  [36].
-Figure 2: Results for aligning LLMs with the Dove preference optimization objective.
-We compare the win-rate against the gold responses of the supervised finetuned (SFT), DPO-aligned and Dove-aligned LLM on the (a) TL;DR summarization and (b) the Anthropic-Helpful datasets.
-In our experiments, we utilize ChatGPT to compare the model responses with the gold responses.
-We generate model responses for three sampling temperatures.
-The results are averaged over three runs of the preference optimization objectives.
+Finally, we report the win-rate of the model responses as the evaluation metric for $500$ unseen instructions from the test sets.
+In particular, we report the win-rate against the gold responses for the model generated responses averaged across three sampling temperatures $\text{T}\in\{0.001,0.5,1.0\}$.
 6.2 Results
-We compare the performance of the SFT, DPO, and Dove aligned models in Figure 2.
-In particular, we report the win-rate against the gold responses for the model responses generated from various sampling temperatures $\text{T}\in\{0.001,0.5,1.0\}$.
-Dove outperforms SFT model.
-We find that the Dove achieves high win-rates across all sampling temperatures.
-Specifically, we observe that Dove outperforms the SFT model by $29.1\%$ and $18\%$ on the close-ended summarization and open-ended dialogue dataset, respectively, averaged across the sampling temperatures.
-This indicates that Dove can utilize the diverse set of feedback from the conditional and joint preferences to align LLMs.
-Dove outperforms DPO algorithm.
-Further, we aim to understand whether Dove is able to tease out useful feedback signals from the combination of the conditional preferences and joint preferences over instruction-response pairs.
-Surprisingly, we find that Dove outperforms DPO by $5.2\%$ and $3.3\%$ win-rate points on the summarization and helpfulness datasets, respectively.
-In addition, the performance of Dove is better than Dove across all the sampling temperatures.
-This highlights that one can improve the alignment of the LLMs by leveraging novel preference acquisition paths without collecting new instruction-response data.
-Further, our results indicate that Dove is a robust alignment algorithm that can elicit high-quality outputs by learning from diverse ranking-based preferences.
-Figure 3: Win-rate against the gold response in the TL;DR and Anthropic-Helpful datasets averaged over three sampling temperatures.
-We study the impact of the joint preferences over non-identical instructions using Dove.
+We compare the performance of the DPO, KTO, and JPO aligned models in Figure 4(a) and 4(b).
+Interestingly, we find that JPO outperforms DPO by $5.2\%$ and $3.3\%$ win-rate points on the summarization and helpfulness datasets, respectively.
+In addition, the performance of JPO is better than DPO across all the sampling temperatures.
+We observe similar trends in comparison to KTO.
+This highlights that one can align LLMs by leveraging novel preference acquisition paths without collecting new instruction-response data.
+Figure 5: Win-rate against the gold response in the TL;DR averaged over three sampling temperatures. We study the impact of the joint preferences over non-identical instructions using JPO.
+6.3 Extending the Results to AlpacaEval
+Similar to Rafailov et al.
+(2023), we show the usefulness of aligning LLMs using joint preferences via JPO on close-ended (e.g., summarization) and open-ended tasks (e.g., dialogues).
+However, we further evaluate the effectiveness of our method on a broad set of instructions in the AlpacaEval2 leaderboard using the length-controlled win-rate metric Li et al.
+(2023).
+Additional experimental details are provided in Appendix E.
+We present the results in Figure 4(c) where we compare JPO with DPO and KTO.
+We find that the JPO-aligned LLM outperforms DPO-aligned LLM by $1.8$ percentage points on the challenging AlpacaEval2 leaderboard using the length-controlled win-rate metric.
+This indicates that the JPO can utilize the joint preferences and elicit helpful and accurate responses for a broad set of instructions.
+7 Ablations
 Impact of Joint Preferences over Non-Identical Instructions.
-Here, we aim to understand the sole impact of joint preferences acquired over non-identical instructions on the performance of the Dove algorithm.
-To do so, we train Dove algorithm with joint feedback data $\mathcal{D}_{H}$ only.
+Here, we aim to understand the sole impact of joint preferences acquired over non-identical instructions on the performance of the JPO algorithm.
+To do so, we train JPO algorithm with joint feedback data $\mathcal{D}_{H}$ only.
 We present the results averaged across the three sampling temperatures in Figure 5.
-We find that training with joint preferences over non-identical instructions achieves $71.7\%$ and $69\%$ win-rate on the summarization and anthropic-helpful datasets, respectively.
+We find that training with joint preferences over non-identical instructions achieves $71.7\%$ win-rate on the summarization dataset.
 This indicates that it is possible to align LLMs with just joint preferences over instruction-response data without any conditional preferences too.
-Furthermore, this highlights that the feedback paths exposed in our setup are robust and effective for eliciting preferred model responses.
-We also show that the training with joint preferences scales with the amount of feedback data using the Dove algorithm in Appendix §J.
-6 Related Work
-Alignment using Reinforcement Learning.
-Large Language Models (LLMs) showcase impressive few-shot performance [11] which is further improved by instruction tuning followed by alignment with preferences data [50, 31, 15, 66, 40, 53].
-This alignment is usually done by first optimizing for a reward model on a preference data [10], followed by aligning the LLMs distribution to maximize the learned reward model (using RLHF/RLAIF).
-In this work, we show that preference acquisition is a complex phenomenon and elicits joint preferences over instruction-response data.
-Further, we study the interactions between joint preferences and the traditional approach of conditional rankings for preference elicitation.
-Reward Free Policy Alignment.
-[37] introduced Direct Preference Optimization (DPO) that mitigates the instability due to PPO for reward maximization by skipping reward modeling and optimizing directly within the model parameter space.
-[32] adds an explicit penalty term to avoid a reduction in the likelihood of preferred pairs.
-[16] improves over DPO by using human-aware loss based on prospect theory.
-Concurrent to our work, [61] explores the comparison between two responses for non-identical instructions.
-However, they do not uncover new preference acquisition paths, and instead rely on augmenting the existing conditional rankings.
-In addition, we introduce Dove, a novel preference optimization, complementary to existing works, that learns from joint preferences over instruction-response data and achieves good performance on diverse tasks.
-We refer to §B for a detailed related work.
+Furthermore, this highlights that the feedback paths exposed in our setup are robust and effective for alignment.
+Impact of Dataset Size.
+In the main experiments, we demonstrated that JPO can learn effectively from a combination of conditional preferences (i.e., 100% of the conditional rankings) and joint preferences over non-identical instructions (of the same size as the conditional preferences).
+To assess the impact of dataset size, we trained JPO using a 50:50 mix of conditional and joint preferences for the TL;DR dataset, with a fixed total size as that of conditional.
+Our results in Figure 6 show that JPO achieves a win rate of 71.9%, outperforming DPO, which was trained on only the conditional preference dataset of the same size, by $4.2$ percentage points.
+Figure 6: Win-rate against the gold response in the TL;DR dataset averaged over three sampling temperatures. We study the impact of dataset size on JPO.
+Data Scaling.
+We aim to understand the impact of increasing the number of preferences collected jointly over instruction-response pairs, for non-identical instructions, on the win-rate against the reference summaries in the TL;DR summarization dataset using JPO algorithm.
+We present the results in Figure 7 for the sampling temperature of $0.001$.
+We find that the win-rate scales from $42.4\%$ to $71.7\%$ as the size of the dataset increases from $100$ to $9000$ comparisons.
+We also observe that the change in the win-rate is within $1\%$ when the dataset size increases from $4000$ to $9000$.
+This highlights that the performance gains are non-linear with the dataset size.
+In the future, it would be pertinent to explore techniques for selecting a subset of joint preference comparisons that result in maximum performance gains.
+Figure 7: Results for scaling the feedback data size on TL;DR summarization dataset.
+We find that the win-rate improves with the increase in the dataset size using the JPO preference optimization objective.
 8 Conclusion
 In this work, we propose a framework that elicits preferences jointly over instruction-response pairs.
-Further, we find that the joint preference optimization uncovers new paths of human reasoning that remain obscured in the traditional approach.
-Additionally, we propose Dove, a novel preference optimization objective for aligning LLMs.
-In our experiments, we show that it outperforms DPO on summarization and dialogue datasets.
+Further, we find that the joint preference optimization uncovers heuristics of human decision making that remain obscured in the traditional approach.
+Additionally, we propose JPO, a novel preference optimization objective for aligning LLMs.
+In our experiments, we show that it outperforms DPO and KTO on summarization and dialogue datasets.
+JPO also outperforms DPO and KTO on AlpacaEval2.
 We note that the number of joint preferences over instruction-response data scales quadratically with the number of instances in the instruction-response dataset.
@@
 While traditional LLM evaluation has focused on conditional rankings, LLM evaluation through joint rankings would be an important future work.
-Appendix A Limitations
+9 Limitations
 While there are various protocols for feedback acquisition, our work is focused on acquiring rankings on a pair of responses under a fixed context or jointly over instruction-response pairs.
@@
 For instance, conditional or joint rankings do not quantify the strengths or weaknesses for a particular task.
-In addition, [8] show that different forms of feedback data often disagree with each other.
+In addition, Bansal et al.
+(2023) show that different forms of feedback data often disagree with each other.
 This highlights at the complex and multidimensional aspects of human preferences.
@@
 Hence, the preferences in our dataset are not represented of the diverse demographics in the world.
-It is pertinent that the future work should study the impact of the diverse groups on the feedback data behaviours and subsequent LLM alignment [63].
-Appendix B Detailed Related Work
-Alignment using Reinforcement Learning.
-Large Language Models (LLMs) showcase impressive zero-shot and few-shot performance [11] with instruction tuning further improving their downstream performance [50].
-Not all responses for a prompt are preferable due to a variety of reasons, thus aligning them with human preference is now a common step in fine-tuning the pipeline.
-This alignment is usually done by first optimizing for a reward model on preference data [10, 24, 8], followed by aligning the LLMs distribution that maximizes the learned reward model using Reinforcement Learning (RLHF) [40, 31], with optional Divergence penalty [53] to avoid deviating from the reference policy.
-Despite the success of RLHF, collecting human feedback can be challenging, however, [15, 28, 66] observe that LLMs can provide feedback motivating Reinforcement Learning through AI feedback (RLAIF).
-In this work, we show that preference acquisition is a complex phenomenon and elicit joint preferences over instruction-response data.
-Further, we study the interactions between joint preferences and traditional approach of conditional rankings for preference elicitation.
-Reward Free Policy Alignment.
-[37] introduced Direct Preference Optimization (DPO) that mitigates the instability due to PPO for reward maximization by optimizing directly within the model parameter space, hence by-passing the reward modeling step.
-[26] extends this framework where instead of two responses, alignment is done over the list of responses while [25] improves DPO using statistical rejection sampling.
-[2] provides an offset in the DPO objective to increase the margins and [32] suggests adding an explicit penalty term to avoid a reduction in the likelihood of preferred pairs over the DPO training.
-Contrary to our work where compare the joint distributions, [61] proposes RPO that compares the conditional likelihood of a winning response with the losing response of another prompt.
-Beyond DPO, more recently [16] proposed a human-aware loss function-based framework using prospect theory named KTO, and [5] proposes IPO that uses human preferences expressed as pairwise preferences.
-Lastly, [64] uses sequence likelihood calibration to align the model from human preference.
-Here, we introduce Dove, a novel preference optimization, complementary to existing works, that learns from joint preferences over instruction-response data and achieves good performance on diverse tasks.
-Appendix C Agreement Analysis.
-Here, we compare the agreement between the human and AI preference decisions for the summarization and open-ended tasks under the two acquisition protocols.
-To establish the gold chosen response in the human annotations, we compare the preference from the two annotators.
-If the two annotators agree in their response (e.g., Response $R_{1}$ $>$ Response $R_{2}$), then it is considered the gold chosen response.
-If the annotators disagree in their preferences (Response $R_{1}$ $>$ Response $R_{2}$ for annotator 1 and A$<$B for annotator 2) then we consider it as tie.
-If an annotator is indecisive while the other annotator makes a decisive preference (e.g., Response $R_{1}$ $=$ Response $R_{2}$ for annotator 1 and A$>$B for annotator 2), then the gold response is the one for which a decisive feedback has been provided i.e., Response $R_{1}$ In addition, the inter-annotator agreement is computed by comparing the preference from the two annotators on the identical comparison.
-In case of disagreements where one of the annotator chooses ‘equal’, the agreement score is considered as $0.5$.
-We apply similar guidelines to compute human-AI agreement.
-Dataset Ranking Protocol Human-Human Human-AI
-Summarization (TL;DR) Two Responses under Fixed Instruction 69% 63%
-Anthropic-Helpful 70.1% 62%
-Summarization (TL;DR) Joint Instruction and Response Pairs 62% 60%
-Anthropic-Helpful 74.8% 71%
-Average 69% 64%
-Table 2: Agreement analysis between within human annotators and gold human feedback and AI (ChatGPT) feedback.
-We perform the agreement calculations for the two ranking protocols: (a) conditional rankings, and (b) joint preferences over instruction-response pairs where instructions are non-identical.
-In addition, we assess the agreement rates over the two datasets: (a) summarization and (b) Anthropic-helpful dataset.
-We present the agreement results in Table 2.
-We find that the average agreement $69\%$ and $64\%$ between the human-human and human-AI annotators, respectively.
-These agreement scores are close to the annotation agreements presented in prior works [22, 8].
-Interestingly, the agreement scores vary based on the underlying distribution of the instruction-response pairs and the choice of ranking protocol.
-Overall, our results highlight that AI provides high-quality feedback across summarization and open-ended dataset for both the feedback protocols.
-Appendix D Proof for Dove subsuming DPO
-Here we highlight a result that reduces Dove into DPO when the prompts are the same.
-Lemma 1.
-Under the case where $\mathcal{D}_{X}=\{(I_{i},R_{i},I_{i},R_{j})\}$, that is, prompts are the same for preferred and not-preferred prompt generation pairs, $\mathcal{L}_{\textsc{DPO}}(\theta;\mathcal{D}_{C},\beta,p_{\text{ref}})=\mathcal{L}_{\textsc{Dove}}(\theta;\mathcal{D}_{X},\beta,p_{\text{ref}})$, where $\mathcal{D}_{C}=\{(I_{j},R_{j}^{w},R_{j}^{\ell})\}$.
-Proof.
-$\mathcal{L}_{\textsc{Dove}}(\theta;\mathcal{D}_{X},\beta,p_{\text{ref}})$ $=\mathbb{E}_{(I_{j}^{w},R_{j}^{w},I_{j}^{\ell},R_{j}^{\ell})\sim\mathcal{D}_{X}}\left[\log\left(\sigma\left(\beta\log\frac{p_{\theta}(R_{i}^{w},I_{i}^{w})}{p_{\text{ref}}(R_{i}^{w},I_{i}^{w})}-\beta\log\frac{p_{\theta}(R_{j}^{\ell},I_{j}^{\ell})}{p_{\text{ref}}(R_{j}^{\ell},I_{j}^{\ell})}\right)\right)\right]$ (3)
-$=\mathbb{E}_{(I_{j}^{w}R_{j}^{w},I_{j}^{\ell},R_{j}^{\ell})\sim\mathcal{D}_{H}}\left[\log\left(\sigma\left(\beta\log\frac{p_{\theta}(R_{i}^{w}|I_{i}^{w})p_{\theta}(I_{i}^{w})}{p_{\text{ref}}(R_{i}^{w}|I_{i}^{w})p_{\text{ref}}(R_{i}^{w}|I_{i}^{w})}\right.\right.\right.$
-$\quad\left.\left.\left.-\beta\log\frac{p_{\theta}(R_{j}^{\ell}|I_{j}^{\ell})p_{\theta}(I_{j}^{\ell})}{p_{\text{ref}}(R_{j}^{\ell},I_{j}^{\ell})p_{\text{ref}}(I_{j}^{\ell})}\right)\right)\right]$
-$={\mathbb{E}}_{(I_{j},R_{j}^{w},R_{j}^{\ell})\sim\mathcal{D}_{C}}\left[\operatorname{log}\left(\sigma\left(\beta\operatorname{log}\frac{p_{\theta}(R_{j}^{w}|I_{j})}{p_{\text{ref}}(R_{j}^{w}|I_{j})}-\beta\operatorname{log}\frac{p_{\theta}(R_{j}^{\ell}|I_{j})}{p_{\text{ref}}(R_{j}^{\ell}|I_{j})}\right)\right)\right]$ (4)
-$=\mathcal{L}_{\textsc{DPO}}(\theta;\mathcal{D}_{C},\beta,p_{\text{ref}})$ (5)
-The proof follows from applying bayes rule and substituting $I_{j}^{w}=I_{j}^{\ell}=I_{j}$. ∎
-Appendix E Dataset Statistics
+It is pertinent that the future work should study the impact of the diverse groups on the feedback data behaviours and subsequent LLM alignment (Zhao et al., 2023).
+Appendix A Ranking Feedback Acquisition Protocol
+Assume a supervised finetuned language model $p_{\text{sft}}$ that is capable of responding to user instructions (e.g., imperative tasks or questions).
+The goal of alignment is to ensure that the SFT model generates high-quality outputs, preferred by humans.
+To do so, we consider a set of instructions =$\mathcal{I}=\{I_{1},\ldots,I_{n}\}$ where $n$ is the number of instructions.
+Further, we consider a set of responses $\{R_{j}^{1},R_{j}^{2},\ldots,R_{j}^{k}\}$ where $k$ is the number of responses for each of the instruction $I_{j}\in\mathcal{I}$.
+This forms a dataset of instructions and their corresponding responses, $\mathcal{D}=\{(I_{j},{R_{j}^{1},R_{j}^{2},\ldots,R_{j}^{k}})\}$.
+[^4] We will drop the iterator over j when defining the dataset for the ease of notation.
+Next, we acquire conditional ranking-based feedback over the collected instruction-response data.
+Under this feedback acquisition protocol, the annotator selects a chosen and rejected response from $\{R_{j}^{x},R_{j}^{y}\}$ conditioned on the instruction $I_{j}$ where $x,y\in\{1,2,\ldots,k\}$.
+The preference decision by the annotator is based on the perceived quality of the responses along various dimensions such as helpfulness (accuracy), coherence (grammar), and harmlessness (safety).
+Formally, the annotator assigns an instruction-conditioned ranking feedback $c(I_{j},R_{j}^{x},R_{j}^{y})\in\{R_{j}^{x},R_{j}^{y},\text{Equal}\}$ where ‘Equal’ indicates that both responses are perceived equally good or bad.
+If $c(I_{j},R_{j}^{x},R_{j}^{y})=R_{j}^{x}$, this implies that the response $R_{j}^{x}$ is the chosen response while the $R_{j}^{y}$ is the rejected response by the annotator.
+As a result, the ranking protocol creates a conditional pairwise feedback data $\mathcal{D}_{C}=\{(I_{j},R_{j}^{x},R_{j}^{y},c(I_{j},R_{j}^{x},R_{j}^{y}))\}$.
+Next, we apply an alignment algorithm on this data to elicit human-preferred responses from the LLM.
+Appendix B Alignment Algorithms
+Rafailov et al.
+(2023) introduced direct preference optimization (DPO) that can align a language model without utilizing on an external reward model.
+Specifically, DPO requires that feedback data should consist of conditional preferences between a pair of responses for a given instruction.
+Additionally, the algorithm assumes a preference dataset $\mathcal{D}_{C}$ and the reference model $p_{\text{ref}}$ which is usually the supervised finetuned language model $p_{\text{sft}}$.
+Specifically, it aims to train an aligned model $p_{\theta}$ using an optimization objective that upweights the conditional probability of the chosen response $p_{\theta}(R_{j}^{w}|I_{j})$ over the rejected response $p_{\theta}(R_{j}^{\ell}|I_{j})$ where $R_{j}^{w}$ and $R_{j}^{\ell}$ are the chosen and rejected response, respectively.
+Formally, the optimization objective for DPO, $\mathcal{L}_{\textsc{DPO}}(\theta;\mathcal{D}_{C},\beta,p_{\text{ref}})$ minimizes the expectation over $(I_{j},R_{j}^{w},R_{j}^{\ell})\sim\mathcal{D}_{C}$:
+$\mathbb{E}\left[\operatorname{log}\left(\sigma\left(\beta\operatorname{log}\frac{p_{\theta}(R_{j}^{w}|I_{j})}{p_{\text{ref}}(R_{j}^{w}|I_{j})}-\beta\operatorname{log}\frac{p_{\theta}(R_{j}^{\ell}|I_{j})}{p_{\text{ref}}(R_{j}^{\ell}|I_{j})}\right)\right)\right]$ (2)
+where $\sigma$ denotes the sigmoid function and $\beta$ is a hyperparameter. Post-alignment, the model generates high-quality outputs for unseen instructions.
+Appendix C Dataset Statistics
 We present the dataset statistics in Table 4.
@@
 Table 4: Statistics for the train split of the summarization and open-ended dialogue datasets.
-Appendix F ChatGPT Prompts
-We present the ChatGPT for acquiring conditional rankings feedback and joint preferences over instruction-response pairs in Table 4 and Table 5, respectively.
-You are a helpful following assistant whose goal is to select the preferred output for a given instruction.
-Answer the question by printing only a single choice from [”Output (a)”, ”Output (b)”] (without quotes) corresponding to the correct answer with no other text.
-#### Annotation Guideline
-In this task, we will ask you to select the preferred output AI model’s responses to instructions.
-You will read examples, which are composed of the following:
-1. an Instruction we give to the AI system
-2. Output (a), the first output from the AI system
-3. Output (b), the second output from the AI system
-Your task is to decide which response is better for each example. There are several dimensions that you can think along. Consider the following dimensions:
-Accuracy: The output sentence should be factually consistent with the instruction and the input.
-Coherence: The output sentence should be easy to understand and free of grammatical errors when read on its own.
-Harmlessness: The output sentence should not contain derogatory, harmful, or toxic connotations.
-You should answer using only Output (a) or Output (b) depending on which response is better.
-### Instruction:
-${instruction}$
-### Output (a):
-${output_{1}}$
-### Output (b):
-${output_{2}}$
-## Preferred Output:
-Figure 4: GPT-3.5-Turbo API prompt for comparisons on identical instructions
-You are a helpful following assistant whose goal is to select the preferred output for a given instruction.
-Answer the question by printing only a single choice from [”Output (a)”, ”Output (b)”] (without quotes) corresponding to the correct answer with no other text.
-#### Annotation Guideline
-In this task, we will ask you to select the preferred output AI model’s responses to instructions.
-You will read examples, which are composed of the following:
-1. Instruction (a), the first input to the AI system and Output (a), the first output from the AI system
-2. Instruction (b), the second input to the AI system and Output (b), the second output from the AI system
-Your task is to decide which response is better for each example. There are several dimensions that you can think along. Consider the following dimensions:
-Accuracy: The output sentence should be factually consistent with the instruction and the input.
-Coherence: The output sentence should be easy to understand and free of grammatical errors when read on its own.
-Harmlessness: The output sentence should not contain derogatory, harmful, or toxic connotations.
-You should answer using only Output (a) or Output (b) depending on which response is better.
-### Instruction (a):
-${instruction_{1}}$
-### Output (a):
-${output_{1}}$
-### Instruction (b):
-${instruction_{2}}$
-### Output (b):
-${output_{2}}$
-## Preferred Output:
-Figure 5: GPT-3.5-Turbo API prompt for comparisons on non-identical instructions
-Appendix G Human Annotation Platform
-We present the screenshots for the human interface in the Figure 6 (conditional rankings) and Figure 7 (joint ranking preferences over instruction-response pairs).
-Figure 6: Human annotation interface for Conditional Rankings
-Figure 7: Human annotation interface for joint preferences over instruction-response pairs.
-Appendix H Qualitative Examples
+Appendix D Proof for JPO subsuming DPO
+We highlight a result that reduces JPO into DPO when the prompts are the same in Lemma E.1.
+Lemma D.1.
+Under the case where $\mathcal{D}_{X}=\{(I_{i},R_{i},I_{i},R_{j})\}$, that is, prompts are the same for preferred and not-preferred prompt generation pairs, $\mathcal{L}_{\textsc{DPO}}(\theta;\mathcal{D}_{C},\beta,p_{\text{ref}})=\mathcal{L}_{\textsc{JPO}}(\theta;\mathcal{D}_{X},\beta,p_{\text{ref}})$, where $\mathcal{D}_{C}=\{(I_{j},R_{j}^{w},R_{j}^{\ell})\}$.
+Proof.
+$\mathcal{L}_{\textsc{JPO}}(\theta;\mathcal{D}_{X},\beta,p_{\text{ref}})$ $=\mathbb{E}_{(I_{j}^{w},R_{j}^{w},I_{j}^{\ell},R_{j}^{\ell})\sim\mathcal{D}_{X}}\left[\log\left(\sigma\left(\beta\log\frac{p_{\theta}(R_{i}^{w},I_{i}^{w})}{p_{\text{ref}}(R_{i}^{w},I_{i}^{w})}-\beta\log\frac{p_{\theta}(R_{j}^{\ell},I_{j}^{\ell})}{p_{\text{ref}}(R_{j}^{\ell},I_{j}^{\ell})}\right)\right)\right]$ (3)
+$=\mathbb{E}_{(I_{j}^{w}R_{j}^{w},I_{j}^{\ell},R_{j}^{\ell})\sim\mathcal{D}_{X}}\left[\log\left(\sigma\left(\beta\log\frac{p_{\theta}(R_{i}^{w}|I_{i}^{w})p_{\theta}(I_{i}^{w})}{p_{\text{ref}}(R_{i}^{w}|I_{i}^{w})p_{\text{ref}}(I_{i}^{w})}-\beta\log\frac{p_{\theta}(R_{j}^{\ell}|I_{j}^{\ell})p_{\theta}(I_{j}^{\ell})}{p_{\text{ref}}(R_{j}^{\ell},I_{j}^{\ell})p_{\text{ref}}(I_{j}^{\ell})}\right)\right)\right]$ (4)
+$=\mathbb{E}_{(I_{j},R_{j}^{w},R_{j}^{\ell})\sim\mathcal{D}_{C}}\left[\operatorname{log}\left(\sigma\left(\beta\operatorname{log}\frac{p_{\theta}(R_{j}^{w}|I_{j})}{p_{\text{ref}}(R_{j}^{w}|I_{j})}-\beta\operatorname{log}\frac{p_{\theta}(R_{j}^{\ell}|I_{j})}{p_{\text{ref}}(R_{j}^{\ell}|I_{j})}\right)\right)\right]$ (5)
+$=\mathcal{L}_{\textsc{DPO}}(\theta;\mathcal{D}_{C},\beta,p_{\text{ref}})$ (6)
+The proof follows from applying bayes rule and substituting $I_{j}^{w}=I_{j}^{\ell}=I_{j}$. ∎
+Appendix E JPO on AlpacaEval2 Leaderboard
+We train Mistral-7B base model on the UltraChat-200K dataset Ding et al.
+(2023) to get the SFT (reference) model.
+Subsequently, we utilize the conditional preference dataset, Ultrafeedback-binarized (60K instances) Cui et al.
+(2023) to align the SFT model using DPO as the baseline algorithm.
+Specifically, we utilize the training setup highlighted in the alignment handbook for SFT and DPO Tunstall et al.
+(2023).
+Since JPO algorithm allows access to joint preferences, we construct non-identical instruction-response tuples by pairing a chosen instruction-response ($I_{chosen},R_{chosen}$) with a rejected instruction-response ($I_{reject},R_{reject}$) from the Ultrafeedback dataset.
+For simplicity, we do not collect new joint preferences for this experiment, and rather utilize the pairings between chosen and rejected instruction-response pairs as a proxy for true joint preference distribution.
+In particular, we train with JPO algorithm for one epoch, and sweep over three learning rates {1e-7, 3e-7, 5e-7} and set the $\beta={0.01}$.
+Post-training, we sample responses from the SFT model, DPO-aligned LLM, KTO-aligned LLM, and JPO-aligned LLM for the instructions in the AlpacaEval2 with a temperature of $0.7$.
+TL;DR Anthropic-Helpful
+Method T = 0.001 T = 0.5 T = 1.0 Average T = 0.001 T = 0.5 T = 1.0 Average
+SFT 46.6 44.9 39.8 43.8 59.1 56.2 56.8 57.4
+DPO Rafailov et al. (2024) 66.5 67.0 69.5 67.7 73.5 72 69.5 71.7
+KTO Ethayarajh et al. (2024) 71.8 71.9 70.6 71.4 72.8 72.9 68.8 71.5
+JPO (Ours) 72.7 71.9 74.2 72.9 76.3 74.5 74.1 75.0
+Table 5: Results for aligning LLMs with the JPO preference optimization objective.
+We compare the win-rate against the gold responses of the supervised finetuned (SFT), DPO-aligned and JPO-aligned LLM on the (a) TL;DR summarization and (b) the Anthropic-Helpful datasets.
+In our experiments, we utilize ChatGPT to compare the model responses with the gold responses.
+We generate model responses for three sampling temperatures.
+The results are averaged over three runs of the preference optimization objectives.
+Appendix F Qualitative Examples
 In this section, we present the qualitative examples to study the interplay between the conditional rankings and the joint preference over instruction-response pairs.
 Here, we acquire ranking feedback from the human annotators and ask them to provide the reasoning for their decision.
-H.1 Anthropic-Helpful Examples
-We present the qualitative examples for the preferences acquired for the Anthropic-helpful dataset in Figure 8, 9, and 10. We present our observations in the figure captions.
-Figure 8: Interplay between the conditional rankings and joint rankings and reasoning acquired from the human annotators for the Anthropic-Helpful dataset.
-In this example, we find that the response B and D are rejected under the conditional rankings.
-When asked to compare the response B and D, humans consider that the response B answers Instruction 1 better than response D answers Instruction 2.
-This indicates that the joint preference humans elicits a decisive feedback between two responses that were rejected under the conditional rankings.
+F.1 Anthropic-Helpful Examples
+We present the qualitative examples for the preferences acquired for the Anthropic-helpful dataset in Figure 8, and 9. We present our observations in the figure captions.
 Figure 8: Interplay between the conditional rankings and joint rankings and reasoning acquired from the human annotators for the Anthropic-Helpful dataset.
@@
 This indicates that a response that was preferred (rejected) under the conditional rankings can still be preferred (rejected) under the joint rankings.
-H.2 TL;DR Summarization Examples
-We present the qualitative examples for the preferences acquired for the TL;DR summarization dataset in Figure 11, 12, and 13. We present our observations in the figure captions.
+F.2 TL;DR Summarization Examples
+We present the qualitative examples for the preferences acquired for the TL;DR summarization dataset in Figure 10, 11, and 12. We present our observations in the figure captions.
 Figure 10: Interplay between the conditional rankings and joint rankings and reasoning acquired from the human annotators for the TL;DR summarization dataset.
@@
 This highlights that a rejected response can be preferred over a non-rejected response under joint rankings.
-Appendix I Alignment Training Details
-I.1 Supervised Finetuning Details
-We present the SFT details in  Table 4. We perform full-finetuning of Mistral-7B using the source code from https://github.com/abacaj/fine-tune-mistral.
+Appendix G Alignment Training Details
+G.1 Supervised Finetuning Details
+We present the SFT details in table  6. We perform full-finetuning of Mistral-7B using the source code from https://github.com/abacaj/fine-tune-mistral.
 Anthropic-Helpful Dataset
@@
 Table 6: Training details for the supervised finetuning of Mistral-7B.
-I.2 Dove
-We present the training details for Dove preference optimization objective in the Table 5.
+G.2 JPO
+We present the training details for JPO preference optimization objective in the Table 7.
 We select the learning rate hyperparameter by sweeping over three learning rates: $\{1e-5,5e-5,5e-4\}$.
-We utilize the TRL library [52] for the DPO source code.
+We utilize the TRL library von Werra et al.
+(2020) for the DPO source code.
 OpenAI TL;DR Summarization Dataset
 Peak Learning Rate 5e-5
-Optimizer AdamW [27]
+Optimizer AdamW (Loshchilov and Hutter, 2017)
 Learning Schedule Cosine
@@
 $\beta$ 0.1
-Table 5: Training details for Dove preference optimization objective. We use the identical settings for DPO.
-Appendix J Trends with Data Scaling
-We aim to understand the impact of increasing the number of preferences collected jointly over instruction-response pairs, for non-identical instructions, on the win-rate against the reference summaries in the TL;DR summarization dataset using Dove algorithm.
-We present the results in Figure 14 for the sampling temperature of $0.001$.
-We find that the win-rate scales from $42.4\%$ to $71.7\%$ as the size of the dataset increases from $100$ to $9000$ comparisons.
-We also observe that the change in the win-rate is within $1\%$ when the dataset size increases from $4000$ to $9000$.
-This highlights that the performance gains are non-linear with the dataset size.
-In the future, it would be pertinent to explore techniques for selecting a subset of joint preference comparisons that result in maximum performance gains.
-Figure 14: Results for scaling the feedback data size on TL;DR summarization dataset.
-We find that the win-rate improves with the increase in the dataset size using the Dove preference optimization objective.
+Table 7: Training details for JPO preference optimization objective. We use the identical settings for DPO.
+Appendix H ChatGPT Prompts
+We present the ChatGPT for acquiring conditional rankings feedback and joint preferences over instruction-response pairs in Table 13 and Table 14, respectively.
+You are a helpful following assistant whose goal is to select the preferred output for a given instruction.
+Answer the question by printing only a single choice from ["Output (a)", "Output (b)"] (without quotes) corresponding to the correct answer with no other text.
+#### Annotation Guideline
+In this task, we will ask you to select the preferred output AI model’s responses to instructions.
+You will read examples, which are composed of the following:
+1. an Instruction we give to the AI system
+2. Output (a), the first output from the AI system
+3. Output (b), the second output from the AI system
+Your task is to decide which response is better for each example. There are several dimensions that you can think along. Consider the following dimensions:
+Accuracy: The output sentence should be factually consistent with the instruction and the input.
+Coherence: The output sentence should be easy to understand and free of grammatical errors when read on its own.
+Harmlessness: The output sentence should not contain derogatory, harmful, or toxic connotations.
+You should answer using only Output (a) or Output (b) depending on which response is better.
+### Instruction:
+${instruction}$
+### Output (a):
+${output_{1}}$
+### Output (b):
+${output_{2}}$
+## Preferred Output:
+Figure 13: GPT-3.5-Turbo API prompt for comparisons on identical instructions
+You are a helpful following assistant whose goal is to select the preferred output for a given instruction.
+Answer the question by printing only a single choice from ["Output (a)", "Output (b)"] (without quotes) corresponding to the correct answer with no other text.
+#### Annotation Guideline
+In this task, we will ask you to select the preferred output AI model’s responses to instructions.
+You will read examples, which are composed of the following:
+1. Instruction (a), the first input to the AI system and Output (a), the first output from the AI system
+2. Instruction (b), the second input to the AI system and Output (b), the second output from the AI system
+Your task is to decide which response is better for each example. There are several dimensions that you can think along. Consider the following dimensions:
+Accuracy: The output sentence should be factually consistent with the instruction and the input.
+Coherence: The output sentence should be easy to understand and free of grammatical errors when read on its own.
+Harmlessness: The output sentence should not contain derogatory, harmful, or toxic connotations.
+You should answer using only Output (a) or Output (b) depending on which response is better.
+### Instruction (a):
+${instruction_{1}}$
+### Output (a):
+${output_{1}}$
+### Instruction (b):
+${instruction_{2}}$
+### Output (b):
+${output_{2}}$
+## Preferred Output:
+Figure 14: GPT-3.5-Turbo API prompt for comparisons on non-identical instructions
+Appendix I Human Annotation Platform
+For human evaluation, we recruit annotators from Amazon Mechanical Turk, and all annotators are fairly paid more than \$18 USD per hour (it varies depending on the time spent on HITs), which is higher than the national minimum wage where the annotators are recruited.
+We present the screenshots for the human interface in the Figure 15 (conditional rankings) and Figure 16 (joint ranking preferences over instruction-response pairs).
+Figure 15: Human annotation interface for Conditional Rankings
+Figure 16: Human annotation interface for joint preferences over instruction-response pairs.